### 28 Dec 2023 v1

# Nicolás Honorato Droguett 11    Kazuhiro Kurita 11 $0$-0002-7638-3322    Tesshu Hanaka 22 $0$-0001-6943-856X    Hirotaka Ono 11 $0$-0003-0845-3947

Abstract
We propose a new model for graph editing problems on intersection graphs. In well-studied graph editing problems, adding and deleting vertices and edges are used as graph editing operations. As a graph editing operation on intersection graphs, we propose moving objects corresponding to vertices. In this paper, we focus on interval graphs as an intersection graph. We give a linear-time algorithm to find the total moving distance for transforming an interval graph into a complete graph. The concept of this algorithm can be applied for (i) transforming a unit square graph into a complete graph over $L_{\infty}$ distance and (ii) attaining the existence of a $k$-clique on unit interval graphs. In addition, we provide LP-formulations to achieve several properties in the associated graph of unit intervals.
Keywords:
Intersection graphs Optimisation Graph edit distance
1 Introduction
The graph editing problem is a problem that consists of editing a graph $G$ so that $G$ satisfies a graph property $\Pi$. Common graph editing operations in well-studied graph editing problems are vertex and edge deletion/addition. The minimum number of such operations required for $G$ to satisfy $\Pi$ is called graph edit distance (GED) [Riesen2015]. When $\Pi$ contains only the target graph, our problem is equivalent to the problem addressed in this paper [Riesen2015]. GED is a metric used to measure the dissimilarity between two given graphs, defined by Sanfeliu and Fu [Sanfeliu1983]. The general problem of determining the GED is known to be NP-hard [Zeng2009]. However, solutions to this problem continue to be designed due to the growing interest in GED as a pairwise graph similarity measure in a wide range of fields such as pattern recognition, image analysis, and molecular analysis [Gao2009, Xiao2013, GarciaHernandez2019]. For instance, in a face recognition scenario, if a facial pattern can be modelled as a graph, and an image is also represented as a graph (called the data graph), then GED can be employed on these graphs to quantify the similarity of the image and the facial model.
In this paper, we propose a new graph edit operation for the graph edit problem on intersection graphs. Given a collection of geometric objects, an intersection graph $G=(V,E)$ is a graph such that $V$ is a collection of geometric objects and $E$ has an edge if two geometric objects intersect. In this graph model, one natural graph edit operation is moving a geometric object. We consider this moving operation as a graph editing operation and address the problem of measuring the similarity between a given graph and a specific graph property. We formulate the problem as follows.
Problem 1.
Given a collection of objects and a graph property $\Pi$, minimise the total moving distance of the objects so that the resulting intersection graph satisfies $\Pi$.
In the above context, we primarily focus on (unit) intervals. This intersection graph is called (unit) interval graph. We identify an interval graph with a collection of intervals. We formally define these graphs in Section 2 and summarise our results in Table 1.
Table 1: Summary of results. In this table, IG, UIG, and USG ($L_{\infty}$) are abbreviations of interval graphs, unit interval graphs, and unit square graphs for $L_{\infty}$ distance, respectively.
Object Property Running time
IG complete $O(n)$
UIG edgeless Poly. (LP)
acyclic Poly. (LP)
has a $k$-clique $O(n\log n)$
does not have a $k$-clique Poly. (LP)
$k$-connected Poly. (LP)
USG ($L_{\infty}$) complete $O(n)$
[…]
As a type of problem that minimises the maximum moving distance, Fomin et al. [fomin2023] addressed the problem of disk dispersal, where a family $\mathcal{S}$ of $n$ disks, an integer $k\geq 0$ and a real $d\geq 0$ are given, and the goal is to decide whether it is possible to obtain a disjoint arrangement by moving at most $k$ disks by at most $d$ distance each. According to the results of Fiala et al. [Fiala2005], this problem is NP-hard when $d=2$ and $k=n$. On the other hand, it was shown by Fomin et al.[fomin2023] that it is FPT when parameterised by $k+d$ and W[1]-hard when parameterised by $k$ if the movement of disks is restricted to rectilinear directions. While some of the cases presented in these works remain open, the system of distant representatives is continuously addressed in the literature [Dumitrescu2015, biedl2021].
This paper is organised as follows. In Section 2, we formally describe the definitions needed to address the above ideas. In Section 3, over Euclidean distance, we show that Problem 1 is solvable in (i) $O(n)$ time for satisfying the property $\Pi_{\texttt{comp}}$ given an interval graph and (ii) $O(n)$ time for unit square graphs for satisfying $\Pi_{\texttt{comp}}$ under standard word-RAM model. In Section 5, we give LP-formulations for satisfying several fundamental graph properties. Lastly, we present our concluding remarks in Section 6.
2 Preliminaries
In this section, we give general definitions that are used throughout the paper. We refer to the basic terminology of graphs, linear programming, and convex functions described in the textbooks [Diestel:BOOK:2010, cormen2009, boyd2004].
An interval $I$ is a closed line segment of length $\mathrm{len}(I)\in\mathbb{R}^{+}$ on the real line. An interval such that $\mathrm{len}(I)=1$ is called a unit interval. The left endpoint $\ell(I)$ of an interval $I$ is the point that satisfies $\ell(I)\leq y$ for any $y\in I$. Similarly, the right endpoint $r(I)$ of $I$ is the point that satisfies $y\leq r(I)$ for any $y\in I$. The centre of $I$ is the point $(r(l)-\ell(I))/2$ and we denote it as $c(I)$.
The left interval set of a collection of intervals $\mathcal{I}$ is the subcollection of intervals to the “left” of a given point $x$. That is, $L(\mathcal{I},x)=\{I\in\mathcal{I}:\>r(I)<x\}$. Similarly, the right interval set is defined as $R(\mathcal{I},x)=\{I\in\mathcal{I}:\>\ell(I)>x\}$. Moreover, the leftmost endpoint $\ell(\mathcal{I})$ of a collection of intervals $\mathcal{I}$ is defined as $\min_{I\in\mathcal{I}}\ell(I)$. Similarly, the rightmost endpoint $r(\mathcal{I})$ of $\mathcal{I}$ is defined as $\max_{I\in\mathcal{I}}r(I)$. The set of endpoints $E(\mathcal{I})$ of $\mathcal{I}$ is defined as $E(\mathcal{I})=\bigcup_{I\in\mathcal{I}}(\ell(I)\cup r(I))$. The subcollection $\mathcal{I}(i,k)$ be the subcollection $\{I_{i},I_{i+1},\ldots,I_{i+k-1}\}$. The endpoint $x_{k}^{\mathcal{I}}$ is the $k$-th endpoint of $\mathcal{I}$. For an arbitrary point $x$, we define two subsets of intervals $E_{\ell}(\mathcal{I},x)=\left\{I\in\mathcal{I}:\>x_{\ell}(I)=x\right\}$ and $E_{r}(\mathcal{I},x)=\left\{I\in\mathcal{I}:\>x_{r}(I)=x\right\}$ to represent the subcollection of intervals of $\mathcal{I}$ having $x$ as the left and right endpoint, respectively.
A unit square is a regular quadrilateral where each side has length $1$, and its centre is positioned at $p_{i}=(x_{i},y_{i})$.
An edgeless graph is a graph $G=(V,E)$ such that $E=\emptyset$. A complete graph is a graph $G=(V,E)$ such that $E=\binom{V}{2}$. An interval graph is an intersection graph $G=(V,E)$ where the vertex set $V=\{v_{1},\ldots,v_{n}\}$ corresponds to a collection of intervals =$\mathcal{I}=\{I_{1},\ldots,I_{n}\}$. There is an edge $(v_{i},v_{j})\in E$ if and only if $I_{i}\cap I_{j}\neq\emptyset$, for any $1\leq i,j\leq n,\>i\neq j$. An interval graph where for all $I\in\mathcal{I}$, $\mathrm{len}(I)=1$, is called unit interval graph. A unit square graph is an intersection graph $G=(V,E)$ where the vertex set $V=\{v_{1},\ldots,v_{n}\}$ corresponds to an unit square collection $\mathcal{S}=\{S_{1},\ldots,S_{n}\}$. There is an edge $(v_{i},v_{j})\in E$ if and only if $S_{i}\cap S_{j}\neq\emptyset$, for any $1\leq i,j\leq n,\>i\neq j$.
[…]
A linear function $f$ for a given set of real numbers $a_{1},\allowbreak a_{2},\ldots,a_{n}$ and a set of variables $x_{1},x_{2},\ldots,x_{n}$ as $f(x_{1},\ldots,x_{n})=\sum^{n}_{j=1}a_{j}x_{j}$. A linear equality for a real number $b$ and linear function $f$ is an equation of the form $f(x_{1},\ldots,x_{n})=b$. Similarly, a linear inequality are equations of the form $\allowbreak f(x_{1},\ldots,x_{n})\leq b$ and $f(x_{1},\ldots,x_{n})\geq b$. We use linear equalities and inequalities to denote linear constraints. A linear-programming problem is the problem of either minimising or maximising a linear function subject to a finite set of linear constraints. We refer in this paper to a linear program as an instance of a linear programming problem described as:
$\begin{array}[]{llll@{}l}\text{minimise}&f(x_{1},x_{2},\ldots,x_{n})&\\ \text{subject to}&\mathcal{C}\\ \end{array}$ (1)
where $\mathcal{C}$ is a finite collection of linear constraints. We avoid expressly showing linear programs in their standard form to maintain legibility in the present work. The linear function $f(x_{1},x_{2},\ldots,x_{n})$ in (1) is called the objective function. An $n$-vector $\bm{x}$ representing an input of $f$ is called a feasible solution if it satisfies all constraints contained in $\mathcal{C}$. A solution $\bm{x}$ has an objective value $f(\bm{x})$. If the objective value $f(\bm{x})$ is minimum (maximum) among all feasible solutions, then $\bm{x}$ is an optimal solution of the linear program.
A set $C$ is convex if the line segment between two arbitrary points in $C$ lies entirely in $C$. Such a set is called a convex set. A function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is convex if its domain $\mathrm{dom}f$ is a convex set and if for all $x,y\in\mathrm{dom}f$ and $\theta$ such that $0\leq\theta\leq 1$, the inequality $f(\theta x+(1-\theta)y)\leq\theta f(x)+(1-\theta)f(y)$ holds. A function satisfying the above is called a convex function. The graph of a function $f$ is the set $\{(x,f(x))\>|\>x\in\mathrm{dom}f\}$. We mainly refer to the geometric interpretation of the above inequality, which says that if $f$ is convex, then an arbitrary line segment from point $(x,f(x))$ to point $(y,f(y))$ lies above the graph of $f$. The epigraph of a function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is the set $\mathrm{epi}f=\{(x,t)\>|\>x\in\mathrm{dom}f,\>f(x)\leq t\}$. A function is convex if and only if its epigraph is a convex set.
3 Linear-time Algorithms for Satisfying $\Pi_{\texttt{comp}}$
This section gives two liner-time algorithms for satisfying $\Pi_{\texttt{comp}}$ on interval graphs and unit square graphs over Euclidean and $L_{\infty}$ distances, respectively. As a remarkable fact, our algorithm runs even if the given collections of intervals and unit squares are not sorted. We first give a linear-time algorithm for interval graphs.
3.1 Interval Graphs
First, observe that if an interval graph is complete, then all the intervals share at least one point. We call such a point as a gathering point. Furthermore, if the value of a gathering point $x$ is known and the intervals are scattered over the line, then intervals in $L(\mathcal{I},x)$ have to be moved to the right, and intervals $R(\mathcal{I},x)$ have to be moved to the left. We try to find an optimal $x$, which is called an optimal gathering point. We show that $E(\mathcal{I})$ contains at least one optimal gathering point. We define a function to calculate the moving distance for an arbitrary interval. In particular, the moving distance function is defined as:
Definition 1.
The moving distance of an interval $I\in\mathcal{I}$ to a point $x$ is a function $d_{I}:\mathbb{R}\rightarrow\mathbb{R}$ defined as:
$d_{I}(x)=\begin{cases}m_{I}(x-c(I)-\frac{\mathrm{len}(I)}{2}),\quad x-\frac{\mathrm{len}(I)}{2}-c(I)>0\\ m_{I}(c(I)-x-\frac{\mathrm{len}(I)}{2}),\quad x+\frac{\mathrm{len}(I)}{2}-c(I)<0\\ 0,\quad\text{otherwise,}\end{cases}$ (2)
where $m_{I}\in\mathbb{R}^{+}$ is an arbitrary constant describing the slope on both sides. The total moving distance of a collection of intervals is then defined as: $D(\mathcal{I},x)=\sum_{I\in\mathcal{I}}d_{I}(x)$.
In simple words, if interval $I$ is moved from the left to $x$, its moving distance is $m_{I}\left(x-c(I)-\mathrm{len}(I)/{2}\right)$; if it is moved to $x$ from the right, its moving distance is the value given by $m_{I}\left(c(I)-x-\mathrm{len}(I)/2\right)$. In all other cases, the interval intersects $x$, and hence its moving distance is 00. For a given interval graph, we say that the moving distance function is unique when $m_{I}=m_{J}$ for any $I,J\in\mathcal{I}$. We state the following property of the above function.
Lemma 1.
The moving distance $d_{I}(x)$ is a convex function.
{toappendix}
Proof.
Notice that $d_{I}(x)$ is a continuous piecewise function. Moreover, $m(x-c(I)-\mathrm{len}(I)/2)$ and $m(c(I)-x-\mathrm{len}(I)/2)$ describe two decreasing and increasing linear functions with zero at $c(I)-\mathrm{len}(I)/2$ and $c(I)+\mathrm{len}(I)/2$, respectively. Hence the epigraph $\mathrm{epi}\>d_{I}$ is the set that lies in the curve of $d_{I}$ (see Fig. 1).
Figure 1: Epigraph of an arbitrary moving distance function $d_{I}(x)$.
We show that $\mathrm{epi}\>d_{I}$ is a convex set. Let $(x_{1},y_{1})$ and $(x_{2},y_{2})$ be two points in $\mathrm{epi}\>d_{I}$. From the definition of $d_{I}$, $y_{1},y_{2}\geq 0$. Without loss of generality, $x_{1}\leq x_{2}$. A point lying on the line segment between $(x_{1},y_{1})$ and $(x_{2},y_{2})$ is represented by $(x^{\prime},y^{\prime})=(x_{1}+\theta\cdot(x_{2}-x_{1}),y_{1}+\theta\cdot(y_{2}-y_{1}))$ for $\theta\in[0,1]$. We consider the following three cases. The first case is $c(I)-\mathrm{len}(I)/2\leq x^{\prime}\leq c(I)+\mathrm{len}(I)/2$. Since $y_{1},y_{2}\geq 0$, $y^{\prime}\geq 0$.
The second case is $x^{\prime}<c(I)-\mathrm{len}(I)/2$. From the definition of $(x^{\prime},y^{\prime})$, we obtain the following formula.
$y^{\prime}$ $\geq-m(c(I)-\mathrm{len}(I)/2-x^{\prime})$
$y_{1}+\theta(y_{2}-y_{1})$ $\geq-m(-x_{1}-\theta(x_{2}-x_{1})+c(I)-\mathrm{len}(I)/2)$
$y_{1}+\theta(y_{2}-y_{1})$ $\geq-m(-x_{1}+c(I)-\mathrm{len}(I)/2)+m\theta(x_{2}-x_{1}).$
Since $y_{1}\geq 0$ and $0\geq-m(-x_{1}+c(I)-\mathrm{len}(I)/2)$, that is, $y_{1}\geq-m(-x_{1}+c(I)-\mathrm{len}(I)/2)$, $(x^{\prime},y^{\prime})$ is contained in $\mathrm{epi}\>d_{I}$ if $\theta(y_{2}-y_{1})\geq m\theta(x_{2}-x_{1})$. We show that $\theta(y_{2}-y_{1})\geq m\theta(x_{2}-x_{1})$ holds. To this end, we further divide the second case into the following three cases.
The first case is the case $x_{2}<c(I)-\mathrm{len}(I)/2$. Since $(x^{\prime},y^{\prime})$ is a point lying on the segment between $(x_{1},y_{1})$ and $(x_{2},y_{2})$, the following inequality holds.
$y_{2}-y_{1}\geq-m(x_{2}-c(I)+\mathrm{len}(I)/2)+m(x_{1}-c(I)+\mathrm{len}(I)/2)=-m(x_{2}-x_{1}).$
Hence $\theta(y_{2}-y_{1})\geq-m\theta(x_{2}-x_{1})$ holds since $\theta\in[0,1]$. Therefore, $(x^{\prime},y^{\prime})$ is contained in $\mathrm{epi}\>d_{I}$
The second case is when $c(I)-\mathrm{len}(I)/2\leq x_{2}\leq c(I)+\mathrm{len}(I)/2$. We have the same bound for $y_{1}\geq 0$ and $y_{2}\geq 0$. Moreover, $20\leq x_{2}-c(I)+\mathrm{len}(I)/$ 2 holds since $x_{2}\geq c(I)-\mathrm{len}(I)/2$. Hence $y_{2}\geq 0\geq-m(x_{2}-c(I)+\mathrm{len}(I)/2)$. As the same discussion, regardless of the value of $x_{1}$, we obtain $y_{1}\geq-m(x_{1}-c(I)+\mathrm{len}(I)/2)$. Therefore, we obtain $y_{2}-y_{1}\geq-m(x_{2}-x_{1})$.
The last case is when $c(I)+\mathrm{len}(I)/2<x_{2}$. In this case, $y_{2}\geq m(x_{2}-c(I)-\mathrm{len}(I)/2)$ and $y_{1}\geq-m(x_{1}-c(I)+\mathrm{len}(I)/2)$ holds. Thus the following inequality holds.
$y_{2}-y_{1}$ $\geq m(x_{2}-c(I)-\mathrm{len}(I)/2)+m(x_{1}-c(I)+\mathrm{len}(I)/2)$
$y_{2}-y_{1}$ $\geq m(x_{1}+x_{2}-2c(I))$
$y_{1}+\theta(y_{2}-y_{1})$ $\geq y_{1}+m\theta(x_{1}+x_{2}-2c(I))$
$y_{1}+\theta(y_{2}-y_{1})$ $\geq-m(x_{1}-c(I)-\mathrm{len}(I)/2)+m\theta(x_{1}+x_{2}-2c(I))$
$y_{1}+\theta(y_{2}-y_{1})$ $\geq-m(x_{1}-c(I)-\mathrm{len}(I)/2-\theta(x_{1}+x_{2})+2\theta c(I)).$
Hence, for proving the equation $\theta(y_{2}-y_{1})\geq m\theta(x_{2}-x_{1})$ holds, we show
$x_{1}-c(I)-\mathrm{len}(I)/2-\theta(x_{1}+x_{2})+2\theta c(I)$ $\leq x_{1}+\theta(x_{2}-x_{1})-c(I)+\mathrm{len}(I)/2$
$\theta c(I)$ $\leq\theta x_{2}+\mathrm{len}(I)/2.$
From the assumption, $c(I)+\mathrm{len}(I)/2<x_{2}$ and the above inequality holds. Consequently, $(x^{\prime},y^{\prime})$ is contained in $\mathrm{epi}\>d_{I}$.
The third case is when $c(I)+\mathrm{len}(I)/2<x^{\prime}$. From the symmetry of the distance function, $(x^{\prime},y^{\prime})$ is contained in $\mathrm{epi}\>d_{I}$ by the same argument of the second case. $\square$
Corollary 1.
The total moving distance function $D(\mathcal{I},x)$ is convex and piecewise linear. In other words, the slope of linear functions defining $D(\mathcal{I},x)$ is non-decreasing.
In what follows, we define a total moving distance function $D(\mathcal{I},x)$ as a sequence of linear functions $D_{1}(\mathcal{I},x),\ldots,D_{2|\mathcal{I}|+1}(\mathcal{I},x)$. For $11\leq i<j\leq 2|\mathcal{I}|+$ 1, the slope of $D_{i}(\mathcal{I},x)$ is smaller than the slope of $D_{j}(\mathcal{I},x)$. Since the total moving distance function is convex, optimal gathering points are located within a range $e_{l}\leq x\leq e_{r}$, where $e_{l},e_{r}\in\mathbb{R}$ and $e_{l}$, $e_{r}$ are also optimal gathering points. In the subsequent paragraphs, $e_{l}$ is the leftmost optimal gathering point and $e_{r}$ is the right most optimal gathering point. More formally, for any optimal gathering point $x$, $e_{l}$ satisfies $e_{l}\leq x$ and $e_{r}$ satisfies $x\leq e_{r}$. The above corollary is used in the algorithm later. Next, we tighten the range of where an optimal gathering point can be found.
Lemma 2.
Let $D(\mathcal{I},x)$ be a total moving distance function defined by linear functions $D_{1}(\mathcal{I},x),\ldots,D_{2|\mathcal{I}|+1}(\mathcal{I},x)$, $\alpha$ be an integer that satisfies the slope of $D_{\alpha}(\mathcal{I},x)$ is non-positive and $D_{\alpha+1}(\mathcal{I},x)$ is non-negative, and $y$ be a real number in the range of $D_{\alpha}(\mathcal{I},x)$. Then, $y$ is an optimal gathering point.
{toappendix}
Proof.
We first show that there exists such $\alpha$. From the definition of a total moving distance function, the slope of $D_{1}(\mathcal{I},x)$ is negative, and the slope of $D_{2|\mathcal{I}|+1}(\mathcal{I},x)$ is positive. Thus such $\alpha$ exists.
We show that $y$ is an optimal gathering point. Let $y^{\prime}$ be a point distinct to $y$. If $y^{\prime}<y$, $D(\mathcal{I},y^{\prime})\geq D(\mathcal{I},y)$ since the slope of functions is non-decrasing by Corollary 1. If $y^{\prime}>y$, then $D(\mathcal{I},y^{\prime})\geq D(\mathcal{I},y)$ since the slope of $D_{\alpha+1}(\mathcal{I},y)$ is non-negative and the slope of functions is non-decrasing by Corollary 1. Therefore, $y$ is an optimal gathering point. $\square$
The above lemma implies the following corollary.
Corollary 2.
There is an optimal gathering point $x$ such that $x$ is an endpoint of an interval in $\mathcal{I}$.
We are now ready to describe a linear-time algorithm for getting an optimal gathering point $x$. Our algorithm is based on a binary search approach. From the convexity of the total moving distance function and Corollary 2, $x$ has to be between the leftmost and rightmost endpoint of $\mathcal{I}$, namely $\ell=\ell(\mathcal{I})$ and $r=r(\mathcal{I})$, respectively. We initially define the point $x$ as the median point of $E(\mathcal{I})$. Next, we inspect $\mathcal{I}$ to obtain $r(L(\mathcal{I},x))$ and $\ell(R(\mathcal{I},x))$, respectively. We then calculate $D(\mathcal{I},x)$, $D(\mathcal{I},r(L(\mathcal{I},x))$ and $D(\mathcal{I},\ell(R(\mathcal{I},x))$ and compare their values. This comparison leads to the procedure described in Algorithm 1.
1 Procedure $\mathcal{I}$)
2       $\ell\leftarrow\ell(\mathcal{I})$ and $r\leftarrow r(\mathcal{I})$. $x$ is the median of $E(\mathcal{I})$. while $D(\mathcal{I},x)>D(\mathcal{I},r(L(\mathcal{I},x))\lor D(\mathcal{I},x)>D(\mathcal{I},\ell(R(\mathcal{I},x))$ do
3             if $D(\mathcal{I},r(L(\mathcal{I},x))<D(\mathcal{I},x)$ then  $r\leftarrow x$ if $D(\mathcal{I},\ell(R(\mathcal{I},x))<D(\mathcal{I},x)$ then  $\ell\leftarrow x$ $x$ is the median of $\{e\in E(\mathcal{I}):\>\ell\leq e\leq r\}$.
4      return $x$
Algorithm 1 A linear-time algorithm for $\Pi=\Pi_{\texttt{comp}}$.
We now show the correctness of this algorithm. We claim that an optimal gathering point is found by repeating the above procedure. To show the correctness of Algorithm 1, we give a characterization of optimal gathering points.
By Corollaries 2 and 2, it is clear that by checking the endpoints of the intervals and its next and previous endpoints, an optimal gathering point can be found. We show that Algorithm 1 finds such a point.
Lemma 3.
Let $x$ be a point finding by Algorithm 1. Let $\ell$ be the endpoint $r(L(\mathcal{I},x))$ and $r$ be the endpoint $\ell(R(\mathcal{I},x))$. Then, $D(\mathcal{I},x)\leq D(\mathcal{I},r)$ and $D(\mathcal{I},x)\leq D(\mathcal{I},\ell)$ hold, that is, $x$ is an optimal gathering point.
{toappendix}
Proof.
From Algorithm 1, $D(\mathcal{I},x)\leq D(\mathcal{I},r)$ and $D(\mathcal{I},x)\leq D(\mathcal{I},\ell)$ hold. Otherwise, Algorithm 1 does not terminate. Moreover, by Corollary 2 and Lemma 2, such $x$ exists and $x$ is an optimal gathering point. $\square$
Lastly, we analyse the time complexity of our algorithm. Algorithm 1 calculates the total moving distance for several points. Straightforwardly, it can be done in $O(n)$ time. To speed up this calculation, we decompose the cost function $D(\mathcal{I},x)$ into three parts $D(L,x)$, $D(R,x)$, and $D(\mathcal{I},x)$, where $L$ and $R$ are a collection of left/right part intervals. Using this decomposition technique, we speed up the calculation of each step.
Theorem 3.1.
Given a collection of intervals, an optimal gathering point can be found in $O(n)$ time.
{toappendix}
Proof.
From Lemmas 2 and 3, Algorithm 1 finds an optimal gathering point $x$. We show that Algorithm 1 runs in $O(n)$ time.
To simplify the analysis, we assume that $\mathcal{I}$ has no two intervals $I$ and $J$ such that $c(I)=c(J)$. If there are such intervals, we identify $I$ and $J$ as one interval $I^{\prime}$ with $m_{I^{\prime}}=m_{I}+m_{J}$. Therefore, this assumption does not lose generality.
In Algorithm 1, we maintain two real numbers $\ell$ and $r$. In what follows, we denote $L(\mathcal{I},\ell)$ and $R(\mathcal{I},r)$ as $L$ and $R$. Assume that we have four values $D(L,\ell)$, $D(R,r)$, $|L|$, and $|R|$. Notice that $D(L,x)$ and $D(R,x)$ can be decomposed into $D(L,\ell)+|L|(x-\ell)$ and $D(R,r)+|R|(r-x)$, respectively. Since $D(\mathcal{I},x)$ can also be decomposed into $D((\mathcal{I}\setminus(L\cup R),x)+D(L,x)+D(R,x)$, we calculate $D(\mathcal{I},x)$ in $O(|(\mathcal{I}\setminus(L\cup R)|)$ time. Moreover, the maintenance of the four values can be done in the same time complexity.
We show that $|\mathcal{I}\setminus(L\cup R)|$ is bounded by the number of endpoints in $\mathcal{I}$ between $\ell$ and $r$. By the definition of $L$ and $R$, both end points of an interval $J\in L\cup R$ are not contained between $\ell$ and $r$. For each interval in $\mathcal{I}\setminus(L\cup R)$, at least one endpoint is contained between $\ell$ and $r$, and $|\mathcal{I}\setminus(L\cup R)|$ is at most the number of endpoints between $\ell$ and $r$. Finally, since $x$ is a median of endpoints in $\mathcal{I}$ between $\ell$ and $r$, and $\mathcal{I}$ has no two intervals with the same centre value, the number of points gets half each time and the total running time is $O(n)$. $\square$
3.1.1 All slopes are the same
We turn our attention to a specific yet sufficiently general case of the above problem setting. In the definition of the moving distance, we allow each interval $I$ to have a distinct $m_{I}$. We consider a case such that, for any $I,J\in\mathcal{I}$, $m_{I}=m_{J}$. In this case, it is not necessary to look for optimal gathering points as they correspond to the $n$- and $n+1$-th endpoints, where $n$ is the number of intervals.
Theorem 3.2.
If the moving distance function of each interval in $\mathcal{I}$ is unique, then the $n$- and $n+1$-th endpoints are optimal gathering points.
{toappendix}
Proof.
Since the moving distance function is unique, the difference of the slopes of the linear functions $D_{1}(\mathcal{I},x),\ldots,D_{2n+1}(\mathcal{I},x)$ is the same. In other words, assume that the slope of $D_{i}(\mathcal{I},x)$ is $s_{i}$. From the definition of the total moving distance function and the fact that the moving distance function is unique, $s_{i+1}-s_{i}=m$ for any $1\leq i\leq 2n$. Moreover, $s_{n+1}=0$ holds since $s_{1}=-mn$. By Lemma 2, the $n$-th point and the $n+1$-th point are optimal gathering points. $\square$
3.2 Unit Square Graphs
We show that Problem 1 applied to unit square graphs is linear-time solvable for $\Pi_{\texttt{comp}}$. Recall that over $L_{\infty}$ distance, the distance between two points $p_{1}=(x_{1},y_{1})$, $p_{2}=(x_{2},y_{2})$ in the plane is: $D(p_{1},p_{2})=\max(|x_{2}-x_{1}|,|y_{2}-y_{1}|)$.
Theorem 3.3.
Given a collection of unit squares $\mathcal{S}$, an optimal gathering point can be found in $O(n)$ time for satisfying $\Pi_{\texttt{comp}}$ over $L_{\infty}$ distance.
{toappendix}
Proof.
Observe that the distance calculation depends on the difference between the $x$- and $y$-axis values individually. Moreover, when employing this specific distance metric, two squares intersect if the distance between their centres is less than 1. Hence it is necessary to ensure that the maximum distance between the $x$- and $y$-axis values remains within 1.
To find an optimal gathering point, we utilise Algorithm 1 and apply it separately to each axis. If $p_{x}$ and $p_{y}$ are the optimal gathering points given by the algorithm applied to $x$- and $y$-axis, respectively, then point $(p_{x},p_{y})$ is an optimal gathering point for this case. This can be done in $O(n)$ time. $\square$
4 Satisfying $\Pi_{k\texttt{-clique}}$ on Unit Interval Graphs in $O(n\log n)$ time
In this section, we address the complexity of the case where a $k$-clique is desired on a given graph. It turns out that this case can be efficiently solvable if the moving distance function is unique. In particular, given a collection of unit intervals $\mathcal{I}$ where the moving distance function is unique, we introduce an $O(n\log n)$-time algorithm to solve problem 1 for satisfying $\Pi_{k\texttt{-clique}}$. We first define the following lemma:
Lemma 4.
Given a collection of unit intervals $\mathcal{I}$ such that it is ordered by the centre of the intervals, an optimal solution for satisfying $\Pi_{k\texttt{-clique}}$ consists of moving intervals $I_{i},\ldots,I_{i+k-1}$ to a gathering point $x$, where $11\leq i\leq n-k+$ 1.
{toappendix}
Proof.
If $\mathcal{I}$ has a $k$-clique, then the statement holds. Thus without loss of generality, assume that $\mathcal{I}$ has no $k$-clique. Let $\mathcal{I}^{\texttt{opt}}$ be an optimal solution. Since $\mathcal{I}^{\texttt{opt}}$ satisfies $\Pi_{k\texttt{-clique}}$, there is a point $x$ such that $\mathcal{I}^{\texttt{opt}}$ has at least $k$ intervals $J_{1}^{\texttt{opt}},\ldots,J_{\ell}^{\texttt{opt}}$ that intersects $x$. If $\ell>k$, then it contradicts the optimality of $\mathcal{I}^{\texttt{opt}}$. Therefore, we assume $\ell=k$.
Let $J_{1},\ldots,J_{k}$ be a collection of intervals in $\mathcal{I}$ that corresponds to intervals $J_{1}^{\texttt{opt}},\ldots,J_{k}^{\texttt{opt}}$. Without loss of generality, for $1\leq i<j\leq|\mathcal{I}|$, the centre of $J_{i}$ is smaller than the centre of $J_{j}$. If $J_{1},\ldots,J_{k}$ is not continuous, we show that it contradicts the optimality of $\mathcal{I}^{\texttt{opt}}$. Notice that $x$ is larger than the centre of $J_{1}$ and smaller than the centre of $J_{k}$ from the optimality of $\mathcal{I}^{\texttt{opt}}$. Let $J^{\prime}$ be an interval in $\mathcal{I}\setminus\{J_{1},\ldots,J_{k}\}$ with the centre larger than $J_{1}$ and smaller than $J_{k}$. Since $J_{1},\ldots,J_{k}$ is not continuous, such $J^{\prime}$ exists. If a point $x$ is left of $J^{\prime}$, then by removing $J_{k}$ and adding $J^{\prime}$, we obtain a solution better than $\mathcal{I}^{\texttt{opt}}$. Moreover, if a point $x$ is right of $J^{\prime}$, then by removing $J_{1}$ and adding $J^{\prime}$, we obtain a better solution. Therefore, $J_{1},\ldots,J_{k}$ is continuous. $\square$
The previous lemma indicates that examining sequences of $k$ consecutive intervals of $\mathcal{I}$ is sufficient to obtain a $k$-clique using the minimum moving distance. Furthermore, using Theorem 3.2, we know that the $k$- and $k+1$-th endpoints are optimal gathering points. Thus we only need to check these endpoints on each sequence to obtain an optimal solution.
The outline of our algorithm is as follows. Let $x_{1}$ and $x_{2}$ be the optimal gathering points of $I(1,k)$ and $I(2,k)$. From Theorem 3.2, $x_{1}$ and $x_{2}$ are medians of $I(1,k)$ and $I(2,k)$. First, we find $x_{1}$ and calculate $D(I(1,k),x_{1})$. Since the difference between $I(1,k)$ and $I(2,k)$ is two, we can find $x_{2}$ and calculate $D(I(2,k),x_{2})$ efficiently. For this purpose, we give formulas for updating the total moving distance.
Lemma 5.
The total moving distance $D(\mathcal{I},x_{n+1}^{\mathcal{I}})$ equals $D(\mathcal{I},x_{n}^{\mathcal{I}})+(x_{n+1}^{\mathcal{I}}-x_{n}^{\mathcal{I}})(|L(\mathcal{I},x_{n+1}^{\mathcal{I}})|-|R(\mathcal{I},x_{n}^{\mathcal{I}})|)$.
{toappendix}
Proof.
The statement holds if $x_{n}^{\mathcal{I}}=x_{n+1}^{\mathcal{I}}$. Thus we assume that $x_{n}^{\mathcal{I}}<x_{n+1}^{\mathcal{I}}$. The set $L(\mathcal{I},x_{n+1}^{\mathcal{I}})$ consists in the intervals contained $L(\mathcal{I},x_{n}^{\mathcal{I}})$ and the intervals having $x_{n}^{\mathcal{I}}$ as right endpoint. That is, $L(\mathcal{I},x_{n+1}^{\mathcal{I}})=L(\mathcal{I},x_{n}^{\mathcal{I}})\cup E_{r}(\mathcal{I},x_{n}^{\mathcal{I}})$. Analogously, $R(\mathcal{I},x_{n+1}^{\mathcal{I}})=R(\mathcal{I},x_{n+1}^{\mathcal{I}})\setminus E_{\ell}(\mathcal{I},x_{n}^{\mathcal{I}})$. Let $\delta$ be $(x_{n+1}^{\mathcal{I}}-x_{n}^{\mathcal{I}})$. For each $I\in L(\mathcal{I},x_{n}^{\mathcal{I}})$, $d_{I}(x_{n+1}^{\mathcal{I}})=d_{I}(x_{n}^{\mathcal{I}})+\delta$. Analogously, for each $I\in R(\mathcal{I},x_{n}^{\mathcal{I}})$, $d_{I}(x_{n+1}^{\mathcal{I}})=d_{I}(x_{n}^{\mathcal{I}})-\delta$. Moreover, the moving distance of $I\in E_{r}(\mathcal{I},x_{n}^{\mathcal{I}})$ to $x_{n+1}^{\mathcal{I}}$ is $d_{I}(x_{n+1}^{\mathcal{I}})=x_{n+1}^{\mathcal{I}}-x_{n}^{\mathcal{I}}$ since $d_{I}(x_{n}^{\mathcal{I}})=0$. Analogously, the moving distance of $I\in E_{\ell}(\mathcal{I},x_{n}^{\mathcal{I}})$ to $x_{n+1}^{\mathcal{I}}$ is $d_{I}(x_{n+1}^{\mathcal{I}})=\delta$ since $d_{I}(x_{n}^{\mathcal{I}})=0$. It follows that $D(\mathcal{I},x_{n+1}^{\mathcal{I}})$ can be reformulated as
$D(\mathcal{I},x_{n+1}^{\mathcal{I}})$ $=\sum_{I\in\mathcal{I}}d_{I}(x_{n+1}^{\mathcal{I}})$
$=\sum_{I\in L(\mathcal{I},x_{n+1}^{\mathcal{I}})}d_{I}(x_{n+1}^{\mathcal{I}})+\sum_{I\in R(\mathcal{I},x_{n+1}^{\mathcal{I}})}d_{I}(x_{n+1}^{\mathcal{I}})$
$=\sum_{I\in L(\mathcal{I},x_{n+1}^{\mathcal{I}})}(d_{I}(x_{n}^{\mathcal{I}})+\delta)+\sum_{I\in R(\mathcal{I},x^{\mathcal{I}}_{n+1})}(d_{I}(x^{\mathcal{I}}_{n})-\delta)$
$=\sum_{I\in L(\mathcal{I},x_{n+1}^{\mathcal{I}})}d_{I}(x^{\mathcal{I}}_{n})+\sum_{I\in R(\mathcal{I},x_{n+1}^{\mathcal{I}})}d_{I}(x^{\mathcal{I}}_{n})+\delta(|L(\mathcal{I},x_{n+1}^{\mathcal{I}})|-|R(\mathcal{I},x_{n}^{\mathcal{I}})|)$
$=D(\mathcal{I},x^{\mathcal{I}}_{n})+(x_{n+1}^{\mathcal{I}}-x_{n}^{\mathcal{I}})(|L(\mathcal{I},x_{n+1}^{\mathcal{I}})|-|R(\mathcal{I},x_{n}^{\mathcal{I}})|).$
Therefore, the formula given in the equation is correct. $\square$
Thus we can calculate the total moving distance efficiently when a gathering point is next to a median. As another update formula, we consider the update formula when one interval is added and one interval is deleted. {toappendix} We first show the following auxiliary lemma. It is used for show the correctness of Lemmas 8 and 7.
Lemma 6.
If =$\mathcal{I}=\{I_{1},\ldots,I_{n}\}$ is an ordered collection of unit intervals by the centre, $I_{n+1}$ is an unit interval such that $c(I_{n})\leq c(I_{n+1})$ and $\mathcal{J}$ is the collection of unit intervals ($(\mathcal{I}\setminus\{I_{1}\})\cup I_{n+1}$, then $\mathcal{I}\cup\{I_{n+1}\}$ has no endpoints between endpoints $x_{n+1}^{\mathcal{I}}$ and $x_{n}^{\mathcal{J}}$. Moreover, $x_{n}^{\mathcal{J}}$ is $\ell(I_{n+1})$ or $x_{n+2}^{\mathcal{I}}$.
Proof.
The only difference between $\mathcal{I}$ and $\mathcal{J}$ is the first and last interval. We show how the indexes change when $I_{1}$ is removed and $I_{n+1}$ is added. We analyse the removal of $I_{1}$ in two cases. The first case is $r(I_{1})\leq x_{n}^{\mathcal{I}}$ and the latter case is $r(I_{1})>x_{n}^{\mathcal{I}}$.
First, if $r(I_{1})\leq x_{n}^{\mathcal{I}}$, then all the endpoints are shifted to the left by two. Thus $x_{n}^{\mathcal{I}}=x_{n-2}^{\mathcal{I}\setminus\{I_{1}\}}$ holds. When $I_{n+1}$ is added, there are further two cases. If $\ell(I_{n+1})<x_{n+1}^{\mathcal{I}}$, then $\ell(I_{n+1})\geq x_{n}^{\mathcal{I}}$. Otherwise, $\ell(I_{n+1})$ would be part of the first $n$ endpoints and thus contained in $\mathcal{I}$. This is a contradiction to the fact that $I_{n+1}$ is contained in $\mathcal{J}$. Consequently, $\ell(I_{n+1})$ is $x_{n-1}^{\mathcal{J}}$ and $x_{n+1}^{\mathcal{I}}$ is $x_{n}^{\mathcal{J}}$. If $\ell(I_{n+1})\geq x_{n+1}^{\mathcal{I}}$, then there are two new points after $x_{n+1}^{\mathcal{I}}$. Hence the closest endpoint to the right of $x_{n+1}^{\mathcal{I}}$ is $x_{n}^{\mathcal{J}}$. In both cases $x_{n+1}^{\mathcal{I}}$ and $x_{n}^{\mathcal{J}}$ are consecutive endpoints.
If $r(I_{1})>x_{n}^{\mathcal{I}}$, then we first show that $r(I_{1})=x_{n+1}^{\mathcal{I}}$. Since the unit interval $I_{1}$ contain no interval $J$ unless $c(I_{1})=c(J)$, every point that contains $I_{1}$ is the left endpoint of an interval in $\mathcal{I}$. Since $r(I_{1})>x_{n}^{\mathcal{I}}$, $I_{i}$ contains $n$ endpoints. Since all $n$ end points are left endpoints, $r(I_{1})=x_{n+1}^{\mathcal{I}}$. When $I_{1}$ is deleted, $x_{n}^{\mathcal{I}}$ becomes $x_{n-1}^{\mathcal{J}}$. If $\ell(I_{n+1})<x_{n+1}^{\mathcal{I}}$, then $\ell(I_{n+1})$ is the closest endpoint to $x_{n+1}^{\mathcal{I}}$ since $\ell(I_{n+1})\geq x_{n}^{\mathcal{I}}$. If $\ell(I_{n+1})\geq x_{n+1}^{\mathcal{I}}$, $\ell(I_{n+1})$ or $x_{n+2}^{\mathcal{I}}=x_{n}^{\mathcal{J}}$ are the closest endpoints to the right of $x_{n+1}^{\mathcal{I}}$. Therefore, in all cases there are no endpoints between $x_{n+1}^{\mathcal{I}}$ and $x_{n}^{\mathcal{J}}$ and $x_{n+1}^{\mathcal{I}}$ is $\ell(I_{n+1})$ or $x_{n+2}^{\mathcal{I}}$. $\square$
1 Procedure $\mathcal{I}$,$k$)
2       Sort the interval set $\mathcal{I}$ by the centre of the intervals. $D_{\texttt{act}}\leftarrow D(\mathcal{I}(1,k),x_{k}^{\mathcal{I}(1,k)})$ Calculates $D(\mathcal{I}(1,k),x_{k+1}^{\mathcal{I}(1,k)})$ using $D_{\texttt{act}}$ $x_{\texttt{act}}=x_{k+1}^{\mathcal{I}(1,k)}$ $D_{\texttt{min}}\leftarrow D_{\texttt{act}},\>\mathcal{I}_{\texttt{min}}\leftarrow\mathcal{I}(1,k),\>x_{\texttt{min}}\leftarrow x_{\texttt{act}}$ while $12\leq i\leq n-k+$ 1 do
3             $x_{\texttt{act}}=x_{k}^{\mathcal{I}(i,k)}$ $D_{\texttt{act}}\leftarrow D(\mathcal{I}(i,k),x_{k}^{\mathcal{I}(i,k)})$ using $D(\mathcal{I}(i-1,k),x_{k+1}^{\mathcal{I}(i-1,k)})$ if $D_{\texttt{act}}<D_{\texttt{min}}$ then  $D_{\texttt{min}}\leftarrow D_{\texttt{act}},\>x_{\texttt{min}}\leftarrow x_{\texttt{act}},\>\mathcal{I}_{\texttt{min}}\leftarrow\mathcal{I}(\left\lceil i/2\right\rceil)$ Calculates $D(\mathcal{I}(i,k),x_{k+1}^{\mathcal{I}(i,k)})$ using $D_{\texttt{act}}$ $i\leftarrow i+1$
4      return ($x_{\texttt{min}},\mathcal{I}_{\texttt{min}})$
Algorithm 2 An $O(n\log n)$-time algorithm for $\Pi=\Pi_{k\texttt{-clique}}$
Lemma 7.
Let $\mathcal{I}$ be a collection of intervals $\{I_{1},\ldots,I_{n}\}$ sorted by the centre, $I_{n+1}$ be an interval such that $c(I_{n})\leq c(I_{n+1})$, and $\mathcal{J}$ be ($(\mathcal{I}\setminus\{I_{1}\})\cup\{I_{n+1}\}$. If $x_{n}^{\mathcal{J}}\geq x_{n+1}^{\mathcal{I}}$ holds, then the total moving distance $D(\mathcal{J},x_{n}^{\mathcal{J}})$ equals $D(\mathcal{I},x_{n+1}^{\mathcal{I}})+|x_{n+1}^{\mathcal{I}}-x_{n}^{\mathcal{J}}|(|L(\mathcal{J},x_{n}^{\mathcal{J}})|-|R(\mathcal{J},x_{n+1}^{\mathcal{I}})|)+d_{I_{n+1}}(x_{n+1}^{\mathcal{I}})-d_{I_{1}}(x_{n+1}^{\mathcal{I}})$.
{toappendix}
Proof.
The statement holds if $x_{n}^{\mathcal{J}}=x_{n+1}^{\mathcal{I}}$. Lemma 6 ensures that $\mathcal{I}\cup\{I_{n+1}\}$ has no points between $x_{n}^{\mathcal{J}}$ and $x_{n+1}^{\mathcal{I}}$. Thus $L(\mathcal{J},x_{n}^{\mathcal{J}})$ equals $(L(\mathcal{I},x_{n+1}^{\mathcal{I}})\cup E_{r}(\mathcal{I},x_{n+1}^{\mathcal{I}}))\setminus\{I_{1}\}$ since $x_{n}^{\mathcal{J}}\geq x_{n+1}^{\mathcal{I}}$. Similarly, the set $R(\mathcal{J},x_{n}^{\mathcal{J}})$ is equal to $(R(\mathcal{I},x_{n+1}^{\mathcal{I}})\setminus E_{\ell}(\mathcal{I},x_{n+1}^{\mathcal{I}}))\cup\{I_{n+1}\}$ if $x_{n}^{\mathcal{J}}\notin I_{n+1}$, otherwise $R(\mathcal{I},x_{n+1}^{\mathcal{I}})\setminus E_{\ell}(\mathcal{I},x_{n+1}^{\mathcal{I}})$. Let $\delta=|x_{n}^{\mathcal{J}}-x_{n+1}^{\mathcal{I}}|$ be the distance between $x_{n}^{\mathcal{J}}$ and $x_{n+1}^{\mathcal{I}}$. The intervals to the left of $x_{n}^{\mathcal{J}}$ in $\mathcal{J}$ are moved further from $x_{n+1}^{\mathcal{I}}$ to the right by $\delta$. In particular, $I\in L(\mathcal{I},x_{n}^{\mathcal{J}})$, $d_{I}(x_{n}^{\mathcal{J}})=d_{I}(x_{n+1}^{\mathcal{I}})+\delta$. Conversely, the intervals in $\mathcal{J}$ to the right of $x_{n}^{\mathcal{J}}$ are moved to $x_{n}^{\mathcal{J}}$ less from $x_{n+1}^{\mathcal{I}}$ to the left by $\delta$. That is, $I\in R(\mathcal{J},x_{n+1}^{\mathcal{I}})$, $d_{I}(x_{n}^{\mathcal{J}})=d_{I}(x_{n+1}^{\mathcal{I}})-\delta$. It follows that the total moving distance can be reformulated as
$D(\mathcal{J},x_{n}^{\mathcal{J}})$ $=\sum_{I\in\mathcal{J}}d_{I}(x_{n}^{\mathcal{J}})$
$=\sum_{I\in L(\mathcal{J},x_{n}^{\mathcal{J}})}d_{I}(x_{n}^{\mathcal{J}})+\sum_{I\in R(\mathcal{J},x_{n}^{\mathcal{J}})}d_{I}(x_{n}^{\mathcal{J}})$
$=\sum_{I\in L(\mathcal{J},x_{n}^{\mathcal{J}})}(d_{I}(x_{n+1}^{\mathcal{I}})+\delta)+\sum_{I\in R(\mathcal{J},x_{n}^{\mathcal{I}})}(d_{I}(x_{n+1}^{\mathcal{I}})-\delta)$
For all $I\in E_{r}(\mathcal{I},x_{n+1}^{\mathcal{I}})$, the moving distance to $x_{n}^{\mathcal{J}}$ equals $d_{I}(x_{n}^{\mathcal{J}})=\delta$ since $d_{I}(x_{n+1}^{\mathcal{I}})=0$. By the definition of $L(\mathcal{J},x_{n}^{\mathcal{J}})$, the sum $\sum_{I\in L(\mathcal{J},x_{n}^{\mathcal{J}})}(d_{I}(x_{n+1}^{\mathcal{I}})+\delta)$ can be reformulated as $\sum_{I\in L(\mathcal{J},x_{n+1}^{\mathcal{I}})}(d_{I}(x_{n+1}^{\mathcal{I}})+\delta)+\delta|E_{r}(\mathcal{I},x_{n+1}^{\mathcal{I}})|$ since $L(\mathcal{J},x_{n+1}^{\mathcal{I}})$ and $E_{r}(\mathcal{I},x_{n+1}^{\mathcal{I}})$ are disjoint.
Secondly, the only difference between $L(\mathcal{J},x_{n+1}^{\mathcal{I}})$ and $L(\mathcal{I},x_{n+1}^{\mathcal{I}})$ is $I_{1}$, thus $\sum_{I\in L(\mathcal{J},x_{n+1}^{\mathcal{I}})}d_{I}(x_{n+1}^{\mathcal{I}})$ is equal to $\sum_{I\in L(\mathcal{I},x_{n+1}^{\mathcal{I}})}d_{I}(x_{n+1}^{\mathcal{I}})-d_{I_{1}}(x_{n+1}^{\mathcal{I}})$. Analogously, $\sum_{I\in R(\mathcal{J},x_{n}^{\mathcal{J}})}d_{I}(x_{n+1}^{\mathcal{I}})$ equals $\sum_{I\in R(\mathcal{I},x_{n+1}^{\mathcal{I}})}(d_{I}(x_{n+1}^{\mathcal{I}})-\delta)-\delta|E_{\ell}(\mathcal{I},x_{n+1}^{\mathcal{I}})|+d_{I_{n+1}}(x_{n+1}^{\mathcal{I}})$. Hence,
$D(\mathcal{J},x_{n}^{\mathcal{J}})=$ $\sum_{I\in L(\mathcal{J},x_{n}^{\mathcal{J}})}(d_{I}(x_{n+1}^{\mathcal{I}})+\delta)+\sum_{I\in R(\mathcal{J},x_{n}^{\mathcal{I}})}(d_{I}(x_{n+1}^{\mathcal{J}})-\delta)$
$=$ $\sum_{I\in L(\mathcal{J},x_{n+1}^{\mathcal{I}})}d_{I}(x_{n+1}^{\mathcal{I}})+\sum_{I\in R(\mathcal{J},x_{n+1}^{\mathcal{I}})}d_{I}(x_{n+1}^{\mathcal{I}})$
$+\delta(|L(\mathcal{J},x_{n}^{\mathcal{I}})|-|R(\mathcal{J},x_{n+1}^{\mathcal{I}})|)+d_{I_{n+1}}(x_{n+1}^{\mathcal{I}})-d_{I_{1}}(x_{n+1}^{\mathcal{I}})$
$=$ $D(\mathcal{I},x_{n+1}^{\mathcal{I}})+\delta(|L(\mathcal{J},x_{n}^{\mathcal{I}})|-|R(\mathcal{J},x_{n+1}^{\mathcal{I}})|)+d_{I_{n+1}}(x_{n+1}^{\mathcal{I}})-d_{I_{1}}(x_{n+1}^{\mathcal{I}}).$
Consequently, the formula given in the equation is correct. $\square$
Lemma 8.
Let =$\mathcal{I}=\{I_{1},\ldots,I_{n}\}$ be an ordered collection of unit intervals by the centre, $I_{n+1}$ be an unit interval such that $c(I_{n})\leq c(I_{n+1})$, $\mathcal{J}$ be the collection of unit intervals ($(\mathcal{I}\setminus\{I_{1}\})\cup I_{n+1}$. If $x_{n}^{\mathcal{J}}<x_{n+1}^{\mathcal{I}}$ holds, then the total moving distance $D(\mathcal{J},x_{n}^{\mathcal{J}})$ equals $D(\mathcal{I},x_{n+1}^{\mathcal{I}})+\delta(|R(\mathcal{J},x_{n}^{\mathcal{J}})|-|L(\mathcal{J},x_{n+1}^{\mathcal{I}})|)+d_{I_{n+1}}(x_{n+1}^{\mathcal{I}})-d_{I_{1}}(x_{n+1}^{\mathcal{I}})$.
{toappendix}
Proof.
Lemma 6 ensures that there are no points between $x_{n}^{\mathcal{J}}$ and $x_{n+1}^{\mathcal{I}}$. Thus $L(\mathcal{J},x_{n}^{\mathcal{J}})$ equals $(L(\mathcal{I},x_{n+1}^{\mathcal{I}})\setminus E_{r}(\mathcal{J},x_{n}^{\mathcal{J}}))\setminus\{I_{1}\}$. Similarly, the set $R(\mathcal{J},x_{n}^{\mathcal{J}})$ is equal to $R(\mathcal{I},x_{n+1}^{\mathcal{I}})\cup E_{\ell}(\mathcal{J},x_{n}^{\mathcal{J}})\cup\{I_{n+1}\}$ if $x_{n}^{\mathcal{J}}\notin I_{n+1}$, otherwise $R(\mathcal{I},x_{n+1}^{\mathcal{I}})\cup E_{\ell}(\mathcal{J},x_{n}^{\mathcal{J}})$. Let $\delta=|x_{n}^{\mathcal{J}}-x_{n+1}^{\mathcal{I}}|$ be the distance between $x_{n}^{\mathcal{J}}$ and $x_{n+1}^{\mathcal{I}}$. The intervals to the left of $x_{n}^{\mathcal{J}}$ in $\mathcal{I}$ are moved less to the right from $x_{n+1}^{\mathcal{I}}$ by $\delta$. In particular, $I\in L(\mathcal{I},x_{n}^{\mathcal{J}})$, $d_{I}(x_{n}^{\mathcal{J}})=d_{I}(x_{n+1}^{\mathcal{I}})-\delta$. Conversely, the intervals in $\mathcal{J}$ to the right of $x_{n}^{\mathcal{J}}$ are moved to $x_{n}^{\mathcal{J}}$ further from $x_{n+1}^{\mathcal{I}}$ to the left by $\delta$. That is, $I\in R(\mathcal{J},x_{n+1}\mathcal{I})$, $d_{I}(x_{n}\mathcal{J})=d_{I}(x_{n+1}^{\mathcal{I}})+\delta$. It follows that the total moving distance can be reformulated as
$D(\mathcal{J},x_{n}^{\mathcal{J}})$ $=\sum_{I\in\mathcal{J}}d_{I}(x_{n}^{\mathcal{J}})$
$=\sum_{I\in L(\mathcal{J},x_{n}^{\mathcal{J}})}d_{I}(x_{n}^{\mathcal{J}})+\sum_{I\in R(\mathcal{J},x_{n}^{\mathcal{J}})}d_{I}(x_{n}^{\mathcal{J}})$
$=\sum_{I\in L(\mathcal{J},x_{n}^{\mathcal{J}})}(d_{I}(x_{n+1}^{\mathcal{I}})-\delta)+\sum_{I\in R(\mathcal{J},x_{n}^{\mathcal{J}})}(d_{I}(x_{n+1}^{\mathcal{I}})+\delta).$
For all $I\in E_{\ell}(\mathcal{J},x_{n+1}^{\mathcal{I}})$, the moving distance to $x_{n}^{\mathcal{J}}$ equals $d_{I}(x_{n}^{\mathcal{J}})=\delta$ since $d_{I}(x_{n+1}^{\mathcal{I}})=0$. By the definition of $L(\mathcal{J},x_{n}^{\mathcal{J}})$, the sum $\sum_{I\in L(\mathcal{J},x_{n}^{\mathcal{I}})}(d_{I}(x_{n+1}^{\mathcal{I}})-\delta)$ can be reformulated as $\sum_{I\in L(\mathcal{J},x_{n+1}^{\mathcal{I}})}(d_{I}(x_{n+1}^{\mathcal{I}})-\delta)-\delta|E_{\ell}(\mathcal{J},x_{n+1}^{\mathcal{I}})|-d_{I_{1}}(x_{n+1}^{\mathcal{I}})$.
Analogously, $\sum_{I\in R(\mathcal{J},x_{n}^{\mathcal{J}})}(d_{I}(x_{n+1}^{\mathcal{I}})+\delta)$ is equal to $\sum_{I\in R(\mathcal{I},x_{n+1}^{\mathcal{I}})}(d_{I}(x_{n+1}^{\mathcal{I}})+\delta)+\delta|E_{\ell}(\mathcal{J},x_{n}^{\mathcal{J}})|+d_{I_{n+1}}(x_{n+1}^{\mathcal{I}})$. Hence,
$D(\mathcal{J},x_{n}^{\mathcal{J}})$ $=\sum_{I\in L(\mathcal{J},x_{n}^{\mathcal{J}})}(d_{I}(x_{n+1}^{\mathcal{I}})-\delta)+\sum_{I\in R(\mathcal{J},x_{n}^{\mathcal{J}})}(d_{I}(x_{n+1}^{\mathcal{I}})+\delta)$
$=\sum_{I\in L(\mathcal{J},x_{n+1}^{\mathcal{I}})}d_{I}(x_{n+1}^{\mathcal{J}})+\sum_{I\in R(\mathcal{J},x_{n+1}^{\mathcal{I}})}d_{I}(x_{n+1}^{\mathcal{I}})$
$+\delta(|R(\mathcal{J},x_{n}^{\mathcal{J}})|-|L(\mathcal{J},x_{n+1}^{\mathcal{I}})|)+d_{I_{n+1}}(x_{n+1}^{\mathcal{I}})-d_{I_{1}}(x_{n+1}^{\mathcal{I}})$
$=D(\mathcal{I},x_{n+1}^{\mathcal{I}})+\delta(|R(\mathcal{J},x_{n}^{\mathcal{J}})|-|L(\mathcal{J},x_{n+1}^{\mathcal{I}})|)+d_{I_{n+1}}(x_{n+1}^{\mathcal{I}})-d_{I_{1}}(x_{n+1}^{\mathcal{I}}).$
It follows that the formula given in the equation is correct. $\square$
Lemmas 8 and 7 give how to update the total moving distance. To update the total moving distance efficiently based on Lemmas 8, 7 and 5, we need to calculate point $x_{j}^{\mathcal{I}(i,k)}$ for any $j$ and the number of left/right intervals of a given point $x$. It can be done using a binary search tree [cormen2009] and we can show the following theorem.
Theorem 4.1.
Given a collection of unit intervals $\mathcal{I}$, if the moving distance function is unique, then an optimal solution can be found in $O(n\log n)$ time.
{toappendix}
Proof.
The correctness of Algorithm 2 can be shown Theorems 3.2, 8 and 7. In each $i$-th loop in line 7, we can find the optimal gathering point of $\mathcal{I}(i,k)$ and calculate $D(\mathcal{I}(i,k),x_{k}^{\mathcal{I}(i,k)})$.
We analyse the time complexity of this algorithm. In the loop of line 7, at each $i$-th iteration it is sufficient to calculate the point $x_{j}^{\mathcal{I}(i,k)}$ for any $j$ and the number of left/right intervals of a given point $x$. These operations can be done in $O(\log n)$ time using a binary search tree [cormen2009]. Since the number of left intervals equals the number of right endpoints of intervals, this calculation can be done in $O(\log n)$ time using a binary search tree. Thus we sort $\mathcal{I}$. The time complexity is $O(n\log n)$. $\square$
{toappendix}
We remark that the above theorem exclusively applies to unit interval graphs where the moving distance function is unique. For interval graphs, we are unable to assume that an optimal solution consists of $k$ consecutive intervals ordered by the centre, as the intervals have arbitrary lengths. For instance, an interval located beyond this consecutive list may intersect with the gathering point if it has sufficient length to reach it. Consequently, Lemma 4 does not hold for interval graphs, and other techniques are required to solve this case. In addition, the restraint applied to the moving distance function allows us to apply Theorem 3.2, otherwise the result of Lemma 4 would also not hold.
5 LP-formulations for Other Fundamental Graph Properties on Unit Interval Graphs
This section deal with the remaining properties, $\Pi_{\texttt{edgeless}}$, $\Pi_{\texttt{acyc}}$, $\overline{\Pi}_{k\texttt{-clique}}$, and $\Pi_{k\texttt{-conn}}$ on unit interval graphs. We show that the problems for satisfying these properties can be formulated using linear programming. It implies that these problems can be solved using polynomial-time algorithms for linear programming.
In the following subsections, we use the absolute value in the objective function to express the formulations. This formulation differs from the formulation in (1), but we assert that a formulation with an absolute value in the objective function can easily be transformed into a linear program. For each variable $x_{i}$ of the objective function in an absolute value, we add two new variables $x_{i}^{\prime}$ and $x_{i}^{\prime\prime}$ and replace $x_{i}$ by $x_{i}^{\prime}-x_{i}^{\prime\prime}$ and $|x_{i}|$ by $x_{i}^{\prime}+x_{i}^{\prime\prime}$, for $1\leq i\leq n$. We also add two constraints $x_{i}^{\prime}\geq 0$ and $x_{i}^{\prime\prime}\geq 0$ to $\mathcal{C}$.
We also add a $\varepsilon$ value to formulations that have strict inequalities, where $\varepsilon\in\mathbb{R^{+}}$ is a sufficiently small number. Strict inequalities lead to formulations where there exists no optimal solution. However, if we use a sufficiently small $\varepsilon$, then we obtain a good upper bound. Thus this modification is reasonable.
5.1 The case $\Pi=\Pi_{\texttt{edgeless}}$
We show that Problem 1 can be formulated as a linear problem for satisfying $\Pi_{\texttt{edgeless}}$. Let =$\mathcal{I}=\{I_{1},\ldots,I_{n}\}$ be a collection of unit intervals such that $c(I_{i})~{}\leq~{}c(I_{i+1})$ for all $11\leq i\leq n-$ 1. We assume that $c(I_{i})\leq c(I_{j})$ for $i<j$. We move each interval such that every pair of intervals has no intersections. Moving an interval may change the order of $\mathcal{I}$. However, the following lemma ensures that the order of $\mathcal{I}$ does not change for every property $\Pi$.
Lemma 9.
Let $D=(d_{1},\ldots,d_{n})$ be a vector describing the distances applied to each interval and $\mathcal{I}^{D}=\{I^{D}_{1},\ldots,I^{D}_{n}\}$ be the collection of intervals such that each centre is $c(I_{i})+d_{i}$. If $\mathcal{I}$ has a pair of intervals $I$ and $J$ satisfying $c(I)\leq c(J)$ and $c(I)+d_{i}>c(J)+d_{j}$, then there is the same collection of intervals such that the total moving distance is at most $\sum_{d_{i}\in D}|d_{i}|$.
{toappendix}
Proof.
Let $I$ and $J$ be two intervals in $\mathcal{I}$. Without loss of generality, assume that $c(I)\leq c(J)$. Suppose that $c(I)+d_{i}\geq c(J)+d_{j}$. Since $c(I)\leq c(J)$, $d_{i}\geq d_{j}$. In such case, by replacing $d_{i}$ and $d_{j}$ with $c(J)-c(I)+d_{j}$ and $c(I)-c(J)+d_{i}$, we obtain $c(I)+(c(J)-c(I)+d_{j})=c(J)+d_{j}$ and $c(J)+(c(I)-c(J)+d_{i})=c(I)+d_{i}$, respectively. Hence the collection of intervals is the same.
We show that $|d_{i}|+|d_{j}|\geq|c(I)-c(J)+d_{i}|+|c(J)-c(I)+d_{j}|$. We consider the following cases. Suppose that $d_{i}\geq c(J)-c(I)$. In this case, $|c(I)-c(J)+d_{i}|=(c(I)-c(J)+d_{i})$. If $c(J)-c(I)+d_{j}\geq 0$, then
$|c(I)-c(J)+d_{i}|+|c(J)-c(I)+d_{j}|=d_{i}+d_{j}\leq|d_{i}|+|d_{j}|.$
If $c(J)-c(I)+d_{j}<0$, then
$|c(I)-c(J)+d_{i}|+|c(J)-c(I)+d_{j}|=d_{i}-d_{j}+2(c(I)-c(J)).$
Since $c(J)\geq c(I)$, $d_{j}<0$. Therefore, $|d_{i}|+|d_{j}|=d_{i}-d_{j}$ and $d_{i}-d_{j}\geq d_{i}-d_{j}+2(c(I)-c(J))$.
Suppose that $d_{i}<c(J)-c(I)$. If $c(J)-c(I)+d_{j}<0$, then
$|c(I)-c(J)+d_{i}|+|c(J)-c(I)+d_{j}|=-d_{i}-d_{j}\leq|d_{i}|+|d_{j}|.$
If $c(J)-c(I)+d_{j}\geq 0$, then
$|c(I)-c(J)+d_{i}|+|c(J)-c(I)+d_{j}|=2(c(J)-c(I))-d_{i}+d_{j}.$
Since $d_{j}\geq c(J)-c(I)$, $|d_{i}|+|d_{j}|=|d_{i}|-d_{j}$. Therefore,
$|d_{i}|-d_{j}+2(c(J)-c(I))-d_{i}+d_{j}$ $=|d_{i}|-d_{i}+2(c(J)-c(I)).$
If $d_{i}\geq 0$, then the statement holds since $|d_{i}|-d_{i}=0$ and $c(J)\geq c(I)$. If $d_{i}<0$, then $-2(d_{i}-(c(J)-c(I)))$. Since $d_{i}<c(J)-c(I)$, $-2(d_{i}-(c(J)-c(I)))<0$ holds. $\square$
Thus $\mathcal{I}$ has no pair of intervals that intersect when the following inequality holds: $c(I_{i+1})-c(I_{i})>1\text{ for }\leq i\leq n-1$. In other words, the distance between $c(I_{i+1})$ and $c(I_{i})$ must be greater than $1$. Observe that the calculation of the distance does not require applying the absolute function, as $c(I_{i+1})\geq c(I_{i})$. It follows that the following linear program with $O(n)$ constraints can be defined:
$\begin{array}[]{llcl@{}l}\text{minimise}&\displaystyle\sum\limits_{i=1}^{n}|x_{i}|&\\ \text{subject to}&(x_{i+1}+c(I_{i+1}))-(x_{i}+c(I_{i}))-\varepsilon\geq 1\\ \end{array}$
Theorem 5.1.
An optimal solution for satisfying $\Pi_{\texttt{edgeless}}$ can be found in polynomial time by linear programming.
{toappendix}
Proof.
We show that the above linear program gives an optimal solution for $\Pi_{\texttt{edgeless}}$. First, given a unit interval graph $(G,\mathcal{I})$, let $D=(d_{1},\ldots,d_{n})$ be a vector describing the distances added to each interval, such that the resulting graph $(G^{D},\mathcal{I}^{D})$ is edgeless. By definition, $\>c(I^{D}_{i})=c(I_{i})+d_{i}$, for all $I^{D}_{i}\in\mathcal{I}^{D}$, $1\leq i\leq n$. We assume that $D$ minimises the total moving distance for $\Pi_{\texttt{edgeless}}$. Since every pair of intervals in $\mathcal{I}^{D}$ does not intersect, the inequality $c(I^{D}_{i+1})-c(I^{D}_{i})=c(I_{i+1})+d_{i+1}-(c(I_{i})+d_{i})\geq 1+\varepsilon$ holds for $11\leq i\leq n-$ 1 and $\varepsilon\in\mathbb{R}^{+}$. In other words, the constraint given in the linear program is satisfied by $D$. Hence $D$ is a feasible solution for the linear program.
In the other direction, let $X=(x_{1},\ldots,x_{n})$ be an output of the linear program for an arbitrary graph $(G,\mathcal{I})$. By Lemma 9, the order of $\mathcal{I}^{X}$ is the same as $\mathcal{I}$. Thus the intersection graph of $\mathcal{I}^{X}$ is edgeless. We assume that $X$ does not minimise the total moving distance. That is, there exists a vector $D=(d_{1},\ldots,d_{n})$ such that it minimises the total moving distance $\sum_{i=1}^{n}|d_{i}|$. Since every pair of intervals in $\mathcal{I}^{D}$ does not intersect, $D$ is a feasible solution for the linear program. Moreover, $\sum_{i=1}^{n}|d_{i}|<\sum_{i=1}^{n}|x_{i}|$, which is a contradiction to the initial assumption. Therefore, $X$ minimises the total moving distance for satisfying $\Pi_{\texttt{edgeless}}$. This concludes the proof. $\square$
5.2 The case $\Pi=\Pi_{\texttt{acyc}}$
We assume =$\mathcal{I}=\{I_{1},\ldots,I_{n}\}$ is sorted by its centre. Moreover, by Lemma 9, we assume that the order of $\mathcal{I}$ does not change. Since an interval graph is chordal [Jungnickel2013, Thm 9.2.8], we can take any 3-tuple from $\mathcal{I}$ of the form $(I_{i-1},I_{i},I_{i+1})$. This tuple does not make a cycle if at least one pair of intervals is separated by a distance greater than 1. Thus the following condition assures an acyclic unit interval graph: $c(I_{i+2})-c(I_{i})>1$ for $21\leq i\leq n-$ 2. The absolute value can be removed. The following linear program with $O(n)$ constraints is defined:
$\begin{array}[]{llcl@{}l}\text{minimise}&\displaystyle\sum\limits_{i=1}^{n}|x_{i}|&\\ \text{subject to}&(x_{i+2}+c(I_{i+2}))-(x_{i}+c(I_{i}))-\varepsilon\geq 1\\ \end{array}$
Theorem 5.2.
An optimal solution for satisfying $\Pi_{\texttt{acyc}}$ can be found in polynomial time by linear programming.
{toappendix}
Proof.
We show that the above formulation gives an optimal solution for $\Pi_{\texttt{acyc}}$. Given a unit interval graph $(G,\mathcal{I})$, let $D=(d_{1},\ldots,d_{n})$ be a vector describing the distances applied to each interval, such that the resulting graph $(G^{D},\mathcal{I}^{D})$ is acyclic. By definition, $c(I^{D}_{i})=c(I_{i})+d_{i}$, for all $I^{D}_{i}\in\mathcal{I}^{D}$, $1\leq i\leq n$. We assume that $D$ minimises the total moving distance. Let $\mathcal{I}^{D}=\{I^{D}_{1},\ldots,I^{D}_{n}\}$ be a collection of intervals that defined by $c(I^{D}_{i})=c(I_{i})+d_{i}$. We denote the intersection graph of $\mathcal{I}^{D}$ as $G^{D}$. Since $G^{D}$ is acyclic, $G^{D}$ has no triangle. Moreover, from Lemma 9, the order of intervals does not change and $c(I^{D}_{i+2})-c(I^{D}_{i})\geq 1+\varepsilon$ holds, for $21\leq i\leq n-$ 2 and $\varepsilon\in\mathbb{R}^{+}$. Thus the constraint given in the linear program is satisfied by $D$. It implies that $D$ is a feasible solution.
In the other direction, let $X=(x_{1},\ldots,x_{n})$ be an output given by the linear program for an arbitrary unit interval graph $(G,\mathcal{I})$. From Lemma 9, the order of intervals does not change. Since $X$ satisfies the constraints, $\mathcal{I}^{X}$ satisfies $\Pi_{\texttt{acyc}}$. We prove that $X$ minimises the total moving distance for satisfying $\Pi_{\texttt{acyc}}$. We assume the contrary, thus there is a vector $D=(d_{1},\ldots,d_{n})$ such that $\sum_{i=1}^{n}|d_{i}|<\sum_{i=1}^{n}|x_{i}|$. The vector $D$ is a feasible solution as it was shown in the previous argument. In other words, the linear program would output $D$ instead of $X$ as $\sum_{i=1}^{n}|d_{i}|<\sum_{i=1}^{n}|x_{i}|$ holds. This is a contradiction to the initial assumption. Therefore, $X$ minimises the total moving distance for $\Pi_{\texttt{acyc}}$. $\square$
5.3 The case $\Pi=\overline{\Pi}_{k\texttt{-clique}}$
We generalise the idea presented for the acyclic case and show a linear program to obtain the minimum moving distance for satisfying $\overline{\Pi}_{k\texttt{-clique}}$.
Lemma 10.
Given a unit interval graph $(G,\mathcal{I})$, a $k$-clique exists in $G$ if and only if $c(I_{i+k-1})-c(I_{i})\leq 1$ for $11\leq i\leq n-k+$ 1.
{toappendix}
Proof.
Suppose that there exists a $k$-clique $K_{k}$ in $G$. This clique $K_{k}$ consists of $k$ vertices, thus $k$ intervals intersect in $\mathcal{I}$. Let $I$ and $J$ be the leftmost and rightmost intervals of the intervals corresponding to the vertices of $K_{k}$, respectively. Moreover, $I$ and $J$ are the $i$-th interval and the $j$-th interval in $\mathcal{I}$. Clearly, $I$ and $J$ intersect, hence $c(J)-c(I)\leq 1$ holds. Since $\mathcal{I}$ is a collection of unit intervals, all intervals between $I$ and $J$ also intersect with $I$ and $J$. Therefore, $j=i+k-1$.
In the other direction, we assume that $c(I_{i+k-1})-c(I_{i})\leq 1$ holds for $11\leq i\leq n-k+$ 1. By the index of both intervals, there are other $k-2$ intervals between $I_{i}$ and $I_{i+k-1}$. Since $c(I_{i+k-1})-c(I_{i})\leq 1$, the distance between all intervals $I_{i},I_{i+1},\ldots,I_{i+k-1}$ is also at most $1$. It follows that this sequence forms a $k$-clique. $\square$
[…]
$\begin{array}[]{llcl@{}l}\text{minimise}&\displaystyle\sum\limits_{i=1}^{n}|x_{i}|&\\ \text{subject to}&(c_{i+k-1}+x_{i+k-1})-(c_{i}+x_{i})-\varepsilon&\geq&1&\\ \end{array}$
Theorem 5.3.
An optimal solution for satisfying $\overline{\Pi}_{k\texttt{-clique}}$ can be found in polynomial time by linear programming.
{toappendix}
Proof.
We show that the above formulation gives an optimal solution for $\overline{\Pi}_{k\texttt{-clique}}$ in a similar way as previous formulations. Given a unit interval graph $(G,\mathcal{I})$, let $D=(d_{1},\ldots,d_{n})$ be a vector describing the distances applied to each interval, such that the resulting graph $(G^{D},\mathcal{I}^{D})$ does not contain a $k$-clique. By definition, $c(I^{D}_{i})=c(I_{i})+d_{i}$, for all $I^{D}_{i}\in\mathcal{I}^{D}$, $1\leq i\leq n$. We assume that $D$ minimises the total moving distance. Since $G^{D}$ does not contain a $k$-clique, $c(I^{D}_{i+k})-c(I^{D}_{i})\geq 1+\varepsilon$ holds, $21\leq i\leq n-$ 2 and $\varepsilon\in\mathbb{R}^{+}$, otherwise the intervals $I^{D}_{i},I^{D}_{i+1},\ldots,I^{D}_{i+k-1}$ form a $k$-clique by Lemma 10. Hence $D$ is a feasible solution.
In the other direction, let $X=(x_{1},\ldots,x_{n})$ be an output given by the linear program for an arbitrary unit interval graph $(G,\mathcal{I})$. From Lemma 9, the order of interval $\mathcal{I}^{X}$ does not change. Since $X$ satisfies the constraints, $\mathcal{I}^{X}$ satisfies $\overline{\Pi}_{k\texttt{-clique}}$. We prove that $X$ minimises the total moving distance for satisfying $\overline{\Pi}_{k\texttt{-clique}}$. We assume the contrary, thus there is a vector $D=(d_{1},\ldots,d_{n})$ such that $\sum_{i=1}^{n}|d_{i}|<\sum_{i=1}^{n}|x_{i}|$. Since $\mathcal{I}^{D}$ has no $k$-cliques and the order of $\mathcal{I}^{D}$ and $\mathcal{I}$ are the same, $D$ is a feasible solution. In other words, the linear program would output $D$ instead of $X$ as $\sum_{i=1}^{n}|d_{i}|<\sum_{i=1}^{n}|x_{i}|$ holds. This is a contradiction to the initial assumption. Therefore, $X$ minimises the total moving distance for $\overline{\Pi}_{k\texttt{-clique}}$. $\square$
5.4 The case $\Pi=\Pi_{k\texttt{-conn}}$
In this section, we show a linear program to solve Problem 1 for satisfying $\Pi_{k\texttt{-conn}}$. We first show the following lemma.
Lemma 11.
A unit interval graph $G$ of $\mathcal{I}$ is $k$-connected if and only if all pairs of the form $I_{i}$, $I_{i+k+1}$ share at least one point, for $1\leq i\leq n-(k+1)$.
{toappendix}
Proof.
Suppose that there is a pair of intervals $I_{i}$ and $I_{i+k+1}$ that do not share a point. Thus by removing $k-1$ intervals $I_{i+1},\ldots,I_{i+k}$, the intersection graph of $\mathcal{I}\setminus\{I_{i+1},\ldots,I_{i+k}\}$ becomes disconnected, that is, $G$ is not $k$-connected.
Suppose that $G$ is not $k$-connected, then there are $k-1$ intervals $I_{x_{1}},\ldots,I_{x_{k}}$ such that the interval graph of =$\mathcal{I}^{\prime}=\mathcal{I}\setminus\{I_{x_{1}},\ldots,I_{x_{k}}\}$ is disconnected. Since $\mathcal{I}^{\prime}$ is disconnected, there are two intervals $I_{y}$ and $I_{z}$ such that for any $y\leq w\leq z$, $I_{w}$ is not contained in $\mathcal{I}^{\prime}$. Since we remove $k-1$ intervals, $z-y\leq k+1$, and $I_{y}$ and $I_{z}$ are desired intervals.
$\square$
Using the above lemma, satisfying $\Pi_{k\texttt{-conn}}$ can be achieved in polynomial time by the following linear program:
$\begin{array}[]{llcl@{}l}\text{minimise}&\displaystyle\sum\limits_{i=1}^{n}|x_{i}|&\\ \text{subject to}&(c_{i+k+1}+x_{i+k+1})-(c_{i}+x_{i})&\leq&1,&\\ \end{array}$
Theorem 5.4.
An optimal solution for satisfying $\Pi_{k\texttt{-conn}}$ can be found in polynomial time by linear programming.
{toappendix}
Proof.
We show that the above formulation gives an optimal solution for $\overline{\Pi}_{k\texttt{-conn}}$ similarly to previous formulations. Given a collection of intervals $\mathcal{I}$, let $D=(d_{1},\ldots,d_{n})$ be a vector describing the distances applied to each interval. The resulting interval is $\mathcal{I}^{D}$ defined by $c(I^{D}_{i})=c(I_{i})+d_{i}$, for all $I^{D}_{i}\in\mathcal{I}^{D}$ and for each $1\leq i\leq n$. The intersection graph of $\mathcal{I}^{D}$ is $\Pi_{k\texttt{-conn}}$. We assume that $D$ is an optimum solution. Since $G^{D}$ is $k$-connected, by Lemma 11 $(c_{i+k+1}+d_{i+k+1})-(c_{i}+d_{i})\leq 1$ holds for each $1\leq i\leq n-(k+1)$. Thus the constraints given in the linear program are satisfied by $D$. It implies that $D$ is a feasible solution.
In the other direction, let $X=(x_{1},\ldots,x_{n})$ be an output given by the linear program for a collection of intervals $\mathcal{I}$. From Lemma 9, the order of interval $\mathcal{I}^{X}$ does not change. Since $X$ satisfies the constraints, $\mathcal{I}^{X}$ satisfies $\Pi_{k\texttt{-conn}}$. We show that $X$ minimises the total moving distance for satisfying $\Pi_{k\texttt{-conn}}$. We assume the contrary, thus there is a vector $D=(d_{1},\ldots,d_{n})$ such that $\sum_{i=1}^{n}|d_{i}|<\sum_{i=1}^{n}|x_{i}|$. Let $\mathcal{I}^{D}$ be a collection of intervals such that each interval $c(I^{D}_{i})=c(I)+d_{i}$. From Lemma 11, $\mathcal{I}^{D}$ satisfies the constraints $(c_{i+k+1}+d_{i+k+1})-(c_{i}+d_{i})\leq 1$ for each $1\leq i\leq n-(k+1)$. Hence $D$ is a feasible solution. In other words, the linear program would output $D$ instead of $X$ as $\sum_{i=1}^{n}|d_{i}|<\sum_{i=1}^{n}|x_{i}|$ holds. This is a contradiction to the assumption. Therefore, $X$ minimises the total moving distance for $\Pi_{k\texttt{-conn}}$. $\square$
6 Concluding Remarks
We have defined an optimisation problem described in Problem 1 and presented several results, as summarised in Table 1. In our present work, the main result is the definition of a new edit model to characterise different scenarios of the presented problem. Several algorithmic results are presented for solving cases of fundamental graph properties.
First, we showed a linear-time algorithm to achieve graph completeness on interval graphs, over Euclidean distance. We demonstrated that graph completeness can be achieved in linear time on unit square graphs over $L_{\infty}$ distance by applying this algorithm. We also depicted the contrast for cases where the moving distance is unique, showing that the optimal gathering points are always the $n$-th and $n+1$-th endpoints in those cases. Using this property, we presented a $O(n\log n)$ running-time algorithm for achieving a $k$-clique on unit interval graphs over Euclidean distance. In addition, we deepen into the property of unit interval graphs used to describe $O(n)$-size linear programs to achieve distinct graph properties on unit interval graphs.

### 09 Jan 2024 v2

@@
 We give a linear-time algorithm to find the total moving distance for transforming an interval graph into a complete graph.
-The concept of this algorithm can be applied for (i) transforming a unit square graph into a complete graph over $L_{\infty}$ distance and (ii) attaining the existence of a $k$-clique on unit interval graphs.
+The concept of this algorithm can be applied for (i) transforming a unit square graph into a complete graph over $L_{1}$ distance and (ii) attaining the existence of a $k$-clique on unit interval graphs.
 In addition, we provide LP-formulations to achieve several properties in the associated graph of unit intervals.
@@
 1 Introduction
-The graph editing problem is a problem that consists of editing a graph $G$ so that $G$ satisfies a graph property $\Pi$.
+The graph editing problem is a problem of editing a graph $G$ so that $G$ satisfies a graph property $\Pi$.
 Common graph editing operations in well-studied graph editing problems are vertex and edge deletion/addition.
 The minimum number of such operations required for $G$ to satisfy $\Pi$ is called graph edit distance (GED) [Riesen2015].
-When $\Pi$ contains only the target graph, our problem is equivalent to the problem addressed in this paper [Riesen2015].
+When $\Pi$ contains only the target graph, our problem is equivalent to the problem addressed in [Riesen2015].
 GED is a metric used to measure the dissimilarity between two given graphs, defined by Sanfeliu and Fu [Sanfeliu1983].
 The general problem of determining the GED is known to be NP-hard [Zeng2009].
-However, solutions to this problem continue to be designed due to the growing interest in GED as a pairwise graph similarity measure in a wide range of fields such as pattern recognition, image analysis, and molecular analysis [Gao2009, Xiao2013, GarciaHernandez2019].
-For instance, in a face recognition scenario, if a facial pattern can be modelled as a graph, and an image is also represented as a graph (called the data graph), then GED can be employed on these graphs to quantify the similarity of the image and the facial model.
+However, solutions to this problem continue to be designed due to the growing interest in GED as a pairwise graph similarity measure in a wide range of fields such as pattern recognition, image analysis, and molecular analysis [Gao2009, GarciaHernandez2019, Xiao2013].
+For instance, in a face recognition scenario, if a facial pattern can be modelled as a graph, and an image is also represented as a graph, then GED can be employed on these graphs to quantify the similarity of the image and the facial model.
 In this paper, we propose a new graph edit operation for the graph edit problem on intersection graphs.
-Given a collection of geometric objects, an intersection graph $G=(V,E)$ is a graph such that $V$ is a collection of geometric objects and $E$ has an edge if two geometric objects intersect.
+Given a collection of geometric objects, an intersection graph $G=(V,E)$ is a graph such that $V$ is a collection of geometric objects represented by vertices and $E$ has an edge if two geometric objects intersect.
 In this graph model, one natural graph edit operation is moving a geometric object.
@@
 We formally define these graphs in Section 2 and summarise our results in Table 1.
-Table 1: Summary of results. In this table, IG, UIG, and USG ($L_{\infty}$) are abbreviations of interval graphs, unit interval graphs, and unit square graphs for $L_{\infty}$ distance, respectively.
+Table 1: Summary of our results. In this table, IG, UIG, and USG ($L_{1}$) are abbreviations of interval graphs, unit interval graphs, and unit square graphs for $L_{1}$ distance, respectively.
 Object Property Running time
@@
 $k$-connected Poly. (LP)
-USG ($L_{\infty}$) complete $O(n)$
+USG ($L_{1}$) complete $O(n)$
 […]
-As a type of problem that minimises the maximum moving distance, Fomin et al.
+As another type of problem that minimises the maximum moving distance, Fomin et al.
 [fomin2023] addressed the problem of disk dispersal, where a family $\mathcal{S}$ of $n$ disks, an integer $k\geq 0$ and a real $d\geq 0$ are given, and the goal is to decide whether it is possible to obtain a disjoint arrangement by moving at most $k$ disks by at most $d$ distance each.
@@
 [Fiala2005], this problem is NP-hard when $d=2$ and $k=n$.
-On the other hand, it was shown by Fomin et al.[fomin2023] that it is FPT when parameterised by $k+d$ and W[1]-hard when parameterised by $k$ if the movement of disks is restricted to rectilinear directions.
-While some of the cases presented in these works remain open, the system of distant representatives is continuously addressed in the literature [Dumitrescu2015, biedl2021].
+On the other hand, it was shown by Fomin et al.[fomin2023] that it is FPT when parameterised by $k+d$ and W[1]-hard when parameterised by $k$ when the movement of disks is restricted to rectilinear directions.
+While some of the cases presented in these works remain open, the system of distant representatives is continuously addressed in the literature [biedl2021, Dumitrescu2015].
 This paper is organised as follows.
 In Section 2, we formally describe the definitions needed to address the above ideas.
-In Section 3, over Euclidean distance, we show that Problem 1 is solvable in (i) $O(n)$ time for satisfying the property $\Pi_{\texttt{comp}}$ given an interval graph and (ii) $O(n)$ time for unit square graphs for satisfying $\Pi_{\texttt{comp}}$ under standard word-RAM model.
+In Section 3, we show that Problem 1 is solvable in (i) $O(n)$ time for satisfying the property $\Pi_{\texttt{comp}}$ given an interval graph over Euclidean distance and (ii) $O(n)$ time for unit square graphs for satisfying $\Pi_{\texttt{comp}}$ over $L_{1}$ distance.
+We use the standard word-RAM model.
+In Section 4, we present an $O(n\log n)$-time algorithm for satisfying $\Pi_{k\texttt{-clique}}$ given a unit interval graph over Euclidean distance.
 In Section 5, we give LP-formulations for satisfying several fundamental graph properties.
@@
 In this section, we give general definitions that are used throughout the paper.
-We refer to the basic terminology of graphs, linear programming, and convex functions described in the textbooks [Diestel:BOOK:2010, cormen2009, boyd2004].
+We refer to the basic terminology of graphs, linear programming, and convex functions described in the textbooks [boyd2004, cormen2009, Diestel:BOOK:2010].
 An interval $I$ is a closed line segment of length $\mathrm{len}(I)\in\mathbb{R}^{+}$ on the real line.
@@
 For an arbitrary point $x$, we define two subsets of intervals $E_{\ell}(\mathcal{I},x)=\left\{I\in\mathcal{I}:\>x_{\ell}(I)=x\right\}$ and $E_{r}(\mathcal{I},x)=\left\{I\in\mathcal{I}:\>x_{r}(I)=x\right\}$ to represent the subcollection of intervals of $\mathcal{I}$ having $x$ as the left and right endpoint, respectively.
-A unit square is a regular quadrilateral where each side has length $1$, and its centre is positioned at $p_{i}=(x_{i},y_{i})$.
+A unit square $S$ is a regular quadrilateral where each side has length $1$, and its centre is positioned at $(s_{x},s_{y})\in\mathbb{R}^{2}$.
 An edgeless graph is a graph $G=(V,E)$ such that $E=\emptyset$.
@@
 3 Linear-time Algorithms for Satisfying $\Pi_{\texttt{comp}}$
-This section gives two liner-time algorithms for satisfying $\Pi_{\texttt{comp}}$ on interval graphs and unit square graphs over Euclidean and $L_{\infty}$ distances, respectively.
-As a remarkable fact, our algorithm runs even if the given collections of intervals and unit squares are not sorted.
+This section presents two liner-time algorithms for satisfying $\Pi_{\texttt{comp}}$ on interval graphs and unit square graphs over Euclidean and $L_{1}$ distances, respectively.
+We remark that our algorithm runs even if the given collections of intervals and unit squares are not sorted.
 We first give a linear-time algorithm for interval graphs.
 3.1 Interval Graphs
-First, observe that if an interval graph is complete, then all the intervals share at least one point.
+Observe that if an interval graph is complete, then all the intervals share at least one point.
 We call such a point as a gathering point.
@@
 We show that Problem 1 applied to unit square graphs is linear-time solvable for $\Pi_{\texttt{comp}}$.
-Recall that over $L_{\infty}$ distance, the distance between two points $p_{1}=(x_{1},y_{1})$, $p_{2}=(x_{2},y_{2})$ in the plane is: $D(p_{1},p_{2})=\max(|x_{2}-x_{1}|,|y_{2}-y_{1}|)$.
+Recall that over $L_{1}$ distance, the distance between two points $p_{1}=(x_{1},y_{1})$, $p_{2}=(x_{2},y_{2})$ in the plane is $|x_{2}-x_{1}|+|y_{2}-y_{1}|$.
+For each unit square $S$, we define the moving distance function $d_{S}$ as follows:
+Definition 2.
+The moving distance of an unit square $S\in\mathcal{S}$ with centre $(s_{x},s_{y})$ to a point $p=(p_{x},p_{y})$ is a function $d_{S}:\mathbb{R}^{2}\rightarrow\mathbb{R}$ defined as:
+$d_{S}(p)=m_{S}(\max(|s_{x}-p_{x}|-1/2,0)+\max(|s_{y}-p_{y}|-1/2,0))$
+where $m_{S}\in\mathbb{R}^{+}$ is an arbitrary constant describing the slope. The total moving distance of a collection of squares is then defined as $D(\mathcal{S},p)=\sum_{S\in\mathcal{S}}d_{S}(p)$.
+For a given collection of unit squares $\mathcal{S}$, we give a linear-time algorithm that finds a point $p$ that minimises the total moving distance function.
+Similarly as of intervals, the point $p$ is an optimal gathering point if $p$ minimises the total moving distance function.
 Theorem 3.3.
-Given a collection of unit squares $\mathcal{S}$, an optimal gathering point can be found in $O(n)$ time for satisfying $\Pi_{\texttt{comp}}$ over $L_{\infty}$ distance.
+Given a collection of unit squares $\mathcal{S}$, an optimal gathering point can be found in $O(n)$ time for satisfying $\Pi_{\texttt{comp}}$ over $L_{1}$ distance.
 {toappendix}
 Proof.
-Observe that the distance calculation depends on the difference between the $x$- and $y$-axis values individually.
-Moreover, when employing this specific distance metric, two squares intersect if the distance between their centres is less than 1.
-Hence it is necessary to ensure that the maximum distance between the $x$- and $y$-axis values remains within 1.
-To find an optimal gathering point, we utilise Algorithm 1 and apply it separately to each axis.
-If $p_{x}$ and $p_{y}$ are the optimal gathering points given by the algorithm applied to $x$- and $y$-axis, respectively, then point $(p_{x},p_{y})$ is an optimal gathering point for this case.
-This can be done in $O(n)$ time.
+Let $\mathcal{S}=\{S_{1},\ldots,S_{n}\}$ be a collection of unit squares.
+We denote the centre of $S_{i}$ as $(s^{i}_{x},s^{i}_{y})$.
+For $\mathcal{S}$, we consider the collection of unit intervals $\mathcal{I}_{x}=\{I_{1},\ldots,I_{n}\}$ such that the centre of each $I_{i}$ is $s^{i}_{x}$.
+Similarly, we define the collection of unit intervals $\mathcal{I}_{y}$.
+Let $p_{x}$ and $p_{y}$ be two optimal gathering points of $\mathcal{I}_{x}$ and $\mathcal{I}_{y}$, respectively.
+We show that the point $p=(p_{x},p_{y})$ minimises the sum of moving distances for $\mathcal{S}$.
+Suppose that $p$ is not an optimal gathering point of $\mathcal{S}$.
+Let $\mathrm{opt}=(\mathrm{opt}_{x},\mathrm{opt}_{y})$ be an optimal gathering point of $\mathcal{S}$.
+Since $D(\mathcal{S},p)>D(\mathcal{S},\mathrm{opt})$, $\sum_{I\in\mathcal{I}_{x}}d_{I}(p_{x})>\sum_{I_{i}\in\mathcal{I}_{x}}d_{I}(\mathrm{opt}_{x})$ or $\sum_{I\in\mathcal{I}_{y}}d_{I}(p_{y})>\sum_{I_{i}\in\mathcal{I}_{y}}d_{I}(\mathrm{opt}_{y})$ hold.
+However, from the definition of $p_{x}$ and $p_{y}$, it contradicts the optimality of $p_{x}$ and $p_{y}$.
 $\square$
@@
 The previous lemma indicates that examining sequences of $k$ consecutive intervals of $\mathcal{I}$ is sufficient to obtain a $k$-clique using the minimum moving distance.
-Furthermore, using Theorem 3.2, we know that the $k$- and $k+1$-th endpoints are optimal gathering points.
+Furthermore, we deduce from Theorem 3.2 that the $k$- and $k+1$-th endpoints are optimal gathering points.
 Thus we only need to check these endpoints on each sequence to obtain an optimal solution.
 The outline of our algorithm is as follows.
-Let $x_{1}$ and $x_{2}$ be the optimal gathering points of $I(1,k)$ and $I(2,k)$.
-From Theorem 3.2, $x_{1}$ and $x_{2}$ are medians of $I(1,k)$ and $I(2,k)$.
-First, we find $x_{1}$ and calculate $D(I(1,k),x_{1})$.
-Since the difference between $I(1,k)$ and $I(2,k)$ is two, we can find $x_{2}$ and calculate $D(I(2,k),x_{2})$ efficiently.
+Let $x_{1}$ and $x_{2}$ be the optimal gathering points of $\mathcal{I}(1,k)$ and $\mathcal{I}(2,k)$.
+From Theorem 3.2, $x_{1}$ and $x_{2}$ are medians of $\mathcal{I}(1,k)$ and $\mathcal{I}(2,k)$.
+First, we find $x_{1}$ and calculate $D(\mathcal{I}(1,k),x_{1})$.
+Since the difference between $\mathcal{I}(1,k)$ and $\mathcal{I}(2,k)$ is two intervals, we can find $x_{2}$ and calculate $D(\mathcal{I}(2,k),x_{2})$ efficiently.
 For this purpose, we give formulas for updating the total moving distance.
@@
 It follows that the formula given in the equation is correct. $\square$
-Lemmas 8 and 7 give how to update the total moving distance.
+Lemmas 8 and 7 provide a method to update the total moving distance.
 To update the total moving distance efficiently based on Lemmas 8, 7 and 5, we need to calculate point $x_{j}^{\mathcal{I}(i,k)}$ for any $j$ and the number of left/right intervals of a given point $x$.
@@
 5 LP-formulations for Other Fundamental Graph Properties on Unit Interval Graphs
-This section deal with the remaining properties, $\Pi_{\texttt{edgeless}}$, $\Pi_{\texttt{acyc}}$, $\overline{\Pi}_{k\texttt{-clique}}$, and $\Pi_{k\texttt{-conn}}$ on unit interval graphs.
+This section addresses the remaining properties, $\Pi_{\texttt{edgeless}}$, $\Pi_{\texttt{acyc}}$, $\overline{\Pi}_{k\texttt{-clique}}$, and $\Pi_{k\texttt{-conn}}$ on unit interval graphs.
 We show that the problems for satisfying these properties can be formulated using linear programming.
-It implies that these problems can be solved using polynomial-time algorithms for linear programming.
+It implies that these problems can be solved by using polynomial-time algorithms for linear programming.
 In the following subsections, we use the absolute value in the objective function to express the formulations.
@@
 Since $c(I)\leq c(J)$, $d_{i}\geq d_{j}$.
-In such case, by replacing $d_{i}$ and $d_{j}$ with $c(J)-c(I)+d_{j}$ and $c(I)-c(J)+d_{i}$, we obtain $c(I)+(c(J)-c(I)+d_{j})=c(J)+d_{j}$ and $c(J)+(c(I)-c(J)+d_{i})=c(I)+d_{i}$, respectively.
+In this case, by replacing $d_{i}$ and $d_{j}$ with $c(J)-c(I)+d_{j}$ and $c(I)-c(J)+d_{i}$, we obtain $c(I)+(c(J)-c(I)+d_{j})=c(J)+d_{j}$ and $c(J)+(c(I)-c(J)+d_{i})=c(I)+d_{i}$, respectively.
 Hence the collection of intervals is the same.
@@
 If $d_{i}\geq 0$, then the statement holds since $|d_{i}|-d_{i}=0$ and $c(J)\geq c(I)$. If $d_{i}<0$, then $-2(d_{i}-(c(J)-c(I)))$. Since $d_{i}<c(J)-c(I)$, $-2(d_{i}-(c(J)-c(I)))<0$ holds. $\square$
-Thus $\mathcal{I}$ has no pair of intervals that intersect when the following inequality holds: $c(I_{i+1})-c(I_{i})>1\text{ for }\leq i\leq n-1$.
-In other words, the distance between $c(I_{i+1})$ and $c(I_{i})$ must be greater than $1$.
-Observe that the calculation of the distance does not require applying the absolute function, as $c(I_{i+1})\geq c(I_{i})$.
+Lemma 9 implies that $\mathcal{I}$ has no pair of intervals that intersect when the following inequality holds: $(c(I_{i+1})+d_{i+1})-(c(I_{i})+d_{i})>1\text{ for }\leq i\leq n-1$.
 It follows that the following linear program with $O(n)$ constraints can be defined:
@@
 We assume =$\mathcal{I}=\{I_{1},\ldots,I_{n}\}$ is sorted by its centre.
-Moreover, by Lemma 9, we assume that the order of $\mathcal{I}$ does not change.
+Moreover, from Lemma 9 we deduce that the order of $\mathcal{I}$ does not change.
 Since an interval graph is chordal [Jungnickel2013, Thm 9.2.8], we can take any 3-tuple from $\mathcal{I}$ of the form $(I_{i-1},I_{i},I_{i+1})$.
-This tuple does not make a cycle if at least one pair of intervals is separated by a distance greater than 1.
-Thus the following condition assures an acyclic unit interval graph: $c(I_{i+2})-c(I_{i})>1$ for $21\leq i\leq n-$ 2.
-The absolute value can be removed.
+This tuple does not make a cycle if at least one pair of intervals does not intersect.
+Thus the condition $c(I_{i+2})-c(I_{i})>1$ for $21\leq i\leq n-$ 2, assures an acyclic unit interval graph.
 The following linear program with $O(n)$ constraints is defined:
@@
 Using the above lemma, satisfying $\Pi_{k\texttt{-conn}}$ can be achieved in polynomial time by the following linear program:
-$\begin{array}[]{llcl@{}l}\text{minimise}&\displaystyle\sum\limits_{i=1}^{n}|x_{i}|&\\ \text{subject to}&(c_{i+k+1}+x_{i+k+1})-(c_{i}+x_{i})&\leq&1,&\\ \end{array}$
+$\begin{array}[]{llcl@{}l}\text{minimise}&\displaystyle\sum\limits_{i=1}^{n}|x_{i}|&\\ \text{subject to}&(c_{i+k+1}+x_{i+k+1})-(c_{i}+x_{i})&\leq&1,&\end{array}$
 Theorem 5.4.
@@
 Several algorithmic results are presented for solving cases of fundamental graph properties.
-First, we showed a linear-time algorithm to achieve graph completeness on interval graphs, over Euclidean distance.
-We demonstrated that graph completeness can be achieved in linear time on unit square graphs over $L_{\infty}$ distance by applying this algorithm.
-We also depicted the contrast for cases where the moving distance is unique, showing that the optimal gathering points are always the $n$-th and $n+1$-th endpoints in those cases.
-Using this property, we presented a $O(n\log n)$ running-time algorithm for achieving a $k$-clique on unit interval graphs over Euclidean distance.
-In addition, we deepen into the property of unit interval graphs used to describe $O(n)$-size linear programs to achieve distinct graph properties on unit interval graphs.
+First, we showed a linear-time algorithm to achieve graph completeness on interval graphs using the minimum moving distance, over Euclidean distance.
+By applying this algorithm, we demonstrated that graph completeness can be achieved in linear time on unit square graphs over $L_{1}$ distance.
+We also depicted the contrast for cases where the moving distance is unique, showing that in those cases the optimal gathering points are always the $n$-th and $n+1$-th endpoints.
+Using this property, we presented a $O(n\log n)$-time algorithm for achieving a $k$-clique on unit interval graphs using the minimum moving distance, over Euclidean distance.
+In addition, we deepen into the properties of unit interval graphs and describe $O(n)$-size linear programs to achieve distinct graph properties on unit interval graphs.